### v1

# Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition

Abstract
Layer removal has emerged as a promising approach for compressing large language models (LLMs) by leveraging redundancy within layers to reduce model size and accelerate inference. However, this technique often compromises internal consistency, leading to performance degradation and instability, with varying impacts across different model architectures. In this work, we propose Taco-SVD, a task-aware framework that retains task-critical singular value directions, preserving internal consistency while enabling efficient compression. Unlike direct layer removal, Taco-SVD preserves task-critical transformations to mitigate performance degradation. By leveraging gradient-based attribution methods, Taco-SVD aligns singular values with downstream task objectives. Extensive evaluations demonstrate that Taco-SVD outperforms existing methods in perplexity and task performance across different architectures while ensuring minimal computational overhead.
[…]
Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition
Kainan Liu1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1Ping An Technology (Shenzhen) Co., Ltd., China 2The Hong Kong University of Science and Technology (Guangzhou) {zhangyong203, chengning211}@pingan.com.cn
footnotetext: † Equal contribution.footnotetext: *Corresponding author.footnotetext: This work was done during Kainan Liu’s internship at Ping An Technology.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023). However, their massive parameter sizes create significant computational and memory challenges, making them difficult to deploy on resource-constrained devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
Figure 1: Zero-shot performance of Taco-SVD and pure layer removal under 20% compression ratio. Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
Despite its simplicity and computational efficiency, pure layer removal introduces two notable challenges, as shown in Figure 1. First, while task accuracy may remain stable in certain cases, perplexity (PPL) often degrades significantly, reflecting disruptions to the model’s internal consistency, which stems from impaired information flow and misaligned intermediate representations (Liang et al., 2024). This increase in PPL suggests that layer removal may affect critical transformations responsible for maintaining coherent representations, potentially reducing the model’s generalization ability on complex or unseen tasks.
Second, the impact of layer removal varies significantly across model architectures. For instance, under the same pruning ratio, LLaMA2 (Touvron et al., 2023) maintains reasonable performance, whereas LLaMA3 (Dubey et al., 2024) experiences a steep performance drop. This observation indicates that the tolerance to layer pruning is architecture-dependent. Such variability underscores a lack of robustness in pure layer removal, where certain task-critical transformations may be unevenly affected, resulting in performance inconsistencies across different architectures.
Our analysis suggests that these issues arise from the removal of transformations that play an essential role in maintaining both internal consistency and task-specific performance. While existing methods, such as LoRA-based fine-tuning (Hu et al., 2022) or lightweight layer replacements (Chen et al., 2024), aim to recover or approximate the lost functionality, they often incur substantial computational costs and do not fully address the challenge of preserving task-critical transformations in a resource-efficient manner.
To address these challenges, we propose Taco-SVD (Task-Aware Singular Value Decomposition for LLM Compression), a novel framework designed to address the limitations of pure layer removal by preserving task-critical transformations. Taco-SVD fundamentally rethinks layer removal by integrating task-aware optimization into singular value decomposition, ensuring that compression decisions align with task-specific objectives. Unlike conventional SVD, which focuses on global matrix reconstruction, Taco-SVD leverages gradient attribution to directly quantify the task-specific importance of singular values. Using a small calibration dataset from a general-purpose, unlabeled source, our method identifies and preserves the most impactful transformations within redundant layers. This approach minimizes the indiscriminate loss of critical task-specific transformations, maintaining internal consistency and representation flow while achieving efficient compression.
To validate our approach, we conducted experiments on tasks such as language generation, commonsense reasoning, and mathematical reasoning using LLaMA and Mistral models. Taco-SVD achieves parameter reduction comparable to pure layer removal while improving perplexity and task performance. Notably, Taco-SVD demonstrates consistent robustness across diverse model architectures, maintaining strong performance without additional training. It also achieves strong performance with minimal calibration data—only 64 samples from a general-purpose, unlabeled dataset, demonstrating its effectiveness. Moreover, fine-tuning further enhances its performance, highlighting Taco-SVD’s robustness and efficiency in compressing large language models.
Figure 2: Overview of Taco-SVD. The figure illustrates how Taco-SVD uses gradient attribution to identify task-critical singular values and applies SVD to compress redundant layers while preserving task-relevant transformations.
2 Related Work
2.1 Layer Removal Techniques and Their Impact on Model Consistency
Layer removal is a structured pruning technique that eliminates redundant layers within Large Language Models (LLMs), aiming to reduce the model’s computational and memory requirements. Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers. SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs. LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
Although effective in reducing model size, layer removal often disrupts internal consistency, leading to degraded task performance and increased perplexity, as shown in our analysis building on the concept introduced by Liang et al. (2024). Additionally, its sensitivity to different architectures limits generalizability, with some models tolerating pruning better than others. These challenges highlight the need for approaches that preserve task-critical information while ensuring robust compression across diverse scenarios.
2.2 Post-Pruning Compensation and Training-Based Recovery Approaches
To mitigate the degradation caused by layer removal, post-pruning compensation methods have been developed. (Kim et al., 2024) introduced Shortened LLaMA, which incorporates LoRA (Hu et al., 2022) to recover the ability of layer removal models. However, LoRA alters the original singular value spectrum, which may diminish the contribution of pre-trained features (Shuttleworth et al., 2024). To address this issue, LLM-Streamline (Chen et al., 2024) trains a lightweight module such as feed-forward neural network (FFN) or a transformer layer to approximate the functionality of the pruned redundant layers. Although effective to some extent, these methods incur significant computational and data costs, limiting their practicality in resource-constrained scenarios.
2.3 Singular Value Decomposition for LLM Compression
Singular Value Decomposition (SVD) is a foundational technique for compressing dense models by decomposing weight matrices into low-rank structures. However, vanilla SVD focuses solely on global matrix reconstruction, providing general approximations without considering task-specific characteristics. Recent methods have enhanced SVD to better align with model and task requirements. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
Despite these advancements, existing methods mainly focus on optimizing compression loss without directly addressing the relationship between singular values and task-specific losses. Taco-SVD bridges this gap by integrating task-aware optimization into SVD, leveraging gradient attribution to retain task-critical singular values and align compressed models with task objectives.
3 Method
3.1 Overview and Motivation
In this paper, we propose a novel task-aware compression framework named Taco-SVD, designed to compress redundant layers in large language models (LLMs) while preserving task-critical transformations. Unlike traditional layer removal methods that directly eliminate redundant components, Taco-SVD integrates task-aware singular value decomposition into the layer removal paradigm. By introducing gradient-based attributions, it systematically aligns the importance of singular values with task-specific objective, ensuring that the most essential components within redundant layers are retained during truncation.
Figure 2 provides an overview of Taco-SVD. The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) applying task-aware singular value decomposition (SVD) to retain the most critical singular values and their corresponding singular vectors. Below, we describe each step in detail.
3.2 Redundant Layer Selection
The first step in Taco-SVD is identifying redundant layers. These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance. Following prior works (Song et al., 2024; Chen et al., 2024), we use cosine similarity to quantify the degree of transformation in a given layer.
For a transformer layer with input hidden state $H_{i}\in\mathbb{R}^{d}$ and output hidden state $H_{i+1}\in\mathbb{R}^{d}$, the cosine similarity is computed as:
$\text{cos}(H_{i},H_{i+1})=\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$ (1)
A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant. Instead of directly removing these layers, Taco-SVD compresses their weight matrices using a task-aware approach to preserve task-relevant transformations.
3.3 Task-Aware Singular Value Decomposition
3.3.1 Motivation for Task-Aware SVD
Singular value decomposition (SVD) is a standard technique for compressing weight matrices by providing a global low-rank approximation. Traditional SVD methods retain the top-$k$ singular values based solely on their magnitudes, which are task-agnostic and often fail to capture task-relevant components. This can lead to the removal of critical transformations necessary for downstream tasks (Hsu et al., 2022; Wang et al., 2024).
To address this limitation, Taco-SVD introduces task awareness into SVD by using gradient-based attribution to evaluate the task-specific importance of each singular value and its corresponding singular vectors. This ensures that only the most task-relevant transformations are preserved.
3.3.2 Singular Value Decomposition (SVD)
For a redundant layer’s weight matrix $W\in\mathbb{R}^{m\times n}$, SVD decomposes $W$ as:
$W=U\Sigma V^{T}$ (2)
where $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{k})$ contains the singular values ($\sigma_{i}\geq 0$), and $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, respectively.
3.3.3 Task-Aware Singular Value Importance
To assess the task-specific importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Task-Aware Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the task loss $\mathcal{L}_{\text{task}}$. Using its first-order Taylor expansion, the alteration in the task loss $L$ caused by $\sigma_{i}$ is approximated as:
$\Delta\mathcal{L}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (3)
where $\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the task loss, and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{task}}$ with respect to $W$ onto the singular directions:
$\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (4)
where $G=\frac{\partial\mathcal{L}_{\text{task}}}{\partial W}$ is the gradient of the task loss with respect to $W$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$. The detailed proof can be found in the Appendix A.1.
Integrating the gradient sensitivity with the magnitude of $\sigma_{i}$, we define $\Phi(\sigma_{i})$ as:
$\Phi(\sigma_{i})=\sigma_{i}\cdot\left(\mathbf{u}_{i}^{T}G\mathbf{v}_{i}\right)$ (5)
where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to task performance, considering both its magnitude and alignment with the task gradient. Larger value of $\Phi(\sigma_{i})$ indicates greater importance.
We rank singular values in descending order of task relevance based on $\Phi(\sigma_{i})$. Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
$W_{\text{task}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (6)
This ensures that $W_{\text{task}}$ preserves the transformations most critical to task performance, achieving efficient compression while maintaining task consistency.
3.4 Layer-wise Low-Rank Decomposition
Taco-SVD applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward. For each redundant layer, weight matrices are compressed using Task-Aware Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$. This ensures a low-rank approximation that preserves task-relevant transformations.
By processing one layer at a time, this approach significantly reduces memory overhead, avoiding the need to load the entire model into memory. Additionally, the backward order ensures that compressing deeper layers does not disrupt the consistency of earlier layers, preserving the overall stability of the model. The detailed implementation for Layer-wise Low-Rank Decomposition is provided in Algorithm 1 in the Appendix A.2.
4 Experiments
4.1 Experimental Setup
To validate the effectiveness of Taco-SVD, we conducted experiments across multiple large language models (LLMs) and datasets, focusing on evaluating task performance, cross-model robustness, and computational efficiency. Below, we detail the models, datasets, baselines, and experimental configurations.
Models
We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral-7b (Jiang et al., 2023). These models enable a comprehensive assessment of our method across varying sizes and architectures, highlighting its generability and robustness.
Datasets
Ten datasets are used as evaluation benchmarks including three language generation datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), C4 (Raffel et al., 2020) and seven datasets for common sense and mathematical reasoning include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018) for common sense reasoning, and MathQA (Amini et al., 2019) for mathematical reasoning.
Metric
We use perplexity and accuracy for language generation tasks and reasoning tasks, respectively. All reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024).
Baselines
We compare Taco-SVD against training-free and training-based compression methods at the same compression ratio. For training-free methods, we select the layer removal technique LaCo (Yang et al., 2024), as well as a structured pruning method SliceGPT (Ashkboos et al., 2024) and the SVD-based LLM compression method ASVD (Yuan et al., 2023). For training-based methods, we compare Taco-SVD with LoRA (Hu et al., 2022) and the layer replacement method LLM-Streamline (Chen et al., 2024) with two variants: LLM-Streamline-FFN and LLM-Streamline-Layer, which replace redundant layers with Feed-Forward Neural Network (FFN) and Transformer Layer (Layer), respectively.
Implementation Details
To ensure a fair comparison, all experimental setups are consistent across all methods. Specifically, for training-free methods, we randomly select 512 samples from the WikiText-2 dataset as calibration data. For training-based methods, all models are fine-tuned on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled. Further experimental details can be found in the Appendix A.3.
Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
LaCo 21.1% 244.53 524.26 243.98 20.2 35.17 57.54 26.23 24.66 52.77 24.29 34.41
SliceGPT 20% 9.04 107.65 38.17 19.2 51.05 59.43 34.84 25.26 60.72 23.25 39.11
ASVD 20% 19.28 171.54 27.93 19.8 55.68 55.96 35.84 24.49 66.21 24.19 40.31
Taco-SVD 20.3% 26.42 77.98 28.88 21.4 51.47 68.19 42.82 35.24 66.59 25.83 44.51
LLaMA2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LaCo 18.1% 39.43 93.98 71.77 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
SliceGPT 21.5% 8.63 99.25 40.7 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
ASVD 20% 9.7 88.39 12.46 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
Taco-SVD 21.6% 16.12 44.09 23.13 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
LaCo 19.5% 27.26 59.14 27.42 27.6 52.25 62.98 42.52 33.36 69.91 25.32 44.85
SliceGPT 21.5% 7.44 99.77 37.65 29.4 58.92 64.8 38.96 31.74 64.42 23.65 44.56
ASVD 20% 8.07 101.1 11.86 28.8 63.55 62.59 46.12 32.68 73.72 24.65 47.44
Taco-SVD 20.6% 11.1 56.57 14.76 26.4 60.62 66.22 46.55 35.07 73.11 24.22 47.46
LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
LaCo 19% 753.98 1577.44 807.12 26 48.78 64.56 33.2 29.61 65.07 30.25 42.50
SliceGPT 20% 52.78 460.73 324.71 14.6 43.48 51.38 29.6 22.78 58.38 21.78 34.57
ASVD 20% 2443.99 4469.54 2842.23 11.8 28.37 48.62 26.63 19.2 54.84 18.79 29.75
Taco-SVD 20.1% 37.86 63.97 42.25 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 1: Zero-shot performance of Taco-SVD and training-free baselines under 20% compression ratio. Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
LoRA 25% 11.56 34.05 14.86 24.6 68.73 70.72 51.18 40.1 71.93 26.37 50.52
LS-FFN 25.2% 16.12 59.98 18.4 21.8 66.88 68.03 49.92 37.8 70.35 27.2 48.85
LS-Layer 24.1% 14.67 54.31 16.61 21.6 68.77 70.09 45.02 39.25 72.25 26.83 49.12
Taco-SVD 24.3% 11.62 44.14 14.85 26.2 69.28 70.32 51.7 40.87 73.07 27.17 51.23
LLaMA-2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LoRA 25% 11.7 35.39 13.95 26.2 57.79 63.06 48.15 33.28 71.16 24.15 46.26
LS-FFN 25.7% 13.94 37.43 14.61 24.6 55.98 62.51 48.32 33.28 71 25.19 45.84
LS-Layer 24% 11.57 32.62 12.91 25.4 59.85 61.88 50.21 35.15 72.91 23.85 47.04
Taco-SVD 24.2% 9.59 30.38 12.22 28.4 60.52 64.56 50.32 36.01 71.76 24.85 48.06
LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
LoRA 25% 8.46 58.15 11.35 26.6 63.05 67.4 53.47 38.4 73.5 25.33 49.68
LS-FFN 25.4% 8.11 76.3 11.22 28.4 61.07 67.17 54.04 38.23 73.29 25.49 49.67
LS-Layer 24.4% 8.23 77.04 11.09 29 62.42 67.25 54.25 39.76 74.21 26.06 50.42
Taco-SVD 24.2% 7.52 38.31 10.42 28.6 63.3 67.96 54.63 38.74 73.88 27.27 50.63
LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
LoRA 22.3% 14.56 21.24 21.8 27.6 69.65 69.38 50.54 41.3 72.69 30.18 51.62
LS-FFN 22% 16.79 24.76 24.14 26.4 70.24 70.17 48.97 40.53 72.03 30.25 51.23
LS-Layer 21.7% 14.27 20.85 20.79 27.2 73.53 69.38 51.22 42.66 73.39 31.56 52.71
Taco-SVD 21% 14.13 20.99 21.73 28.2 72.01 69.93 51.5 43.26 73.5 31.26 52.81
Table 2: Performance comparison of Taco-SVD + FT and training-based baselines under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
4.2 Results
We conducted comprehensive experiments to evaluate Taco-SVD from four key perspectives: (a) zero-shot compression performance on language generation and reasoning tasks, (b) effectiveness as an initialization method compared to training-based approaches, (c) robustness at high compression ratios, and (d) compression efficiency and inference speed on real hardware.
Training-Free Compression
To evaluate the generalizability of Taco-SVD, we compared its performance with training-free baselines under approximately 20% compression ratio across four models with varying scales and architectures: Mistral-7B, LLaMA2-7B, LLaMA2-13B, and LLaMA3.1-8B-Instruct. The results, summarized in Table 1, demonstrate that Taco-SVD consistently achieves strong performance across all models and tasks, maintaining competitive perplexity and accuracy even on challenging architectures like LLaMA3. Notably, Taco-SVD exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse architectures.
Training-based Compression
In this section, we further demonstrate that Taco-SVD can provide better initialization for fine-tuning, enabling faster and more accurate recovery of the compressed model compared to other methods. To ensure that the number of trainable parameters remains approximately the same across all methods, we retain only 10% of the parameters in each redundant layer and set only the parameters within these redundant layers to be trainable. As shown in Table 2, Taco-SVD consistently outperforms LoRA and LLM-Streamline across different LLMs. In addition, we plot the validation loss curves during the training process for each method in the Appendix A.4, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
Performance at High Compression Ratio
To evaluate the performance and stability of our proposed method at high compression ratios, we compare Taco-SVD with the baselines on LLaMA2-7B at compression ratios of 25%, 30%, 35% and 40%. The perplexity (PPL) on WikiText-2, PTB and the average accuracy across all reasoning benchmarks are shown in Figure 3, with detailed results for each benchmark presented in the Appendix A.5. We make the following three observations: (a) Most layer removal methods cannot tolerate high compression ratios, with performance dropping sharply at higher compression levels. (b) Compared to training-free baselines, Taco-SVD consistently achieves better accuracy across all compression ratios. (c) Fine-tuning Taco-SVD significantly recovers the accuracy of compressed models, even at high compression ratios.
Figure 3: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and baselines under 20% to 40% compression ratio. Top: Perplexity on WikiText-2 (left) and PTB (right). Bottom: Average accuracy on seven reasoning benchmarks.
Compression Costs and Acceleration Effect
Taco-SVD can not only compress LLMs at very low cost but also improve inference efficiency on real hardware. To demonstrate this, we measure the compression time and peak memory consumption of Taco-SVD when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. These metrics, summarized in Table 3, indicate that Taco-SVD achieves a good balance between compression time and memory consumption. Additionally, to assess the acceleration effect of Taco-SVD, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by layer removal and Taco-SVD with different sequence lengths and batch sizes. As shown in Figure 4, Taco-SVD consistently ensures an enhancement in the generation speed and achieves an acceleration effect comparable to pure layer removal. It is important to note that although Taco-SVD retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance.
Model
Average
Compression Time(h)
Memory
Consumption(MiB)
LaCO 0.05 45046
ASVD 5.5 17632
SliceGPT 0.6 18195
Taco-SVD 0.16 29050
Table 3: Compression time and peak memory usage of Taco-SVD and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
Figure 4: Throughput of LLaMA2-7B and Taco-SVD compressed model under 25% compression ratio on a single A100 GPU. Top: Throughput across different sequence lengths (batch size = 32). Bottom: Throughput across different batch sizes (sequence length = 32).
4.3 Ablation Study
Calibration Data
We conducted ablation studies to investigate the impact of calibration dataset and the number of calibration data used for singular value gradient attribution. Table 4 summarizes the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio using different calibration datasets and varying numbers of calibration data from WikiText-2. The results demonstrate that Taco-SVD is robust to both the choice of calibration dataset and the quantity of calibration data. Notably, even with as few as 64 calibration samples, Taco-SVD achieves competitive performance, highlighting its efficiency in leveraging minimal data for effective compression. We report the detailed results in the Appendix A.6.
Ablation Type
Calibration
Dataset
WikiText-2 PTB
Average
Accuracy
Varying Dataset WikiText-2 37.86 63.97 47.12
C4 40.54 71.42 46.17
Varying Number 64 46.5 86.51 47.06
128 39.91 76.41 46.93
256 38.73 79.13 46.67
512 37.86 63.97 47.12
Table 4: Comparison of Taco-SVD using different types and amounts of calibration data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
One-shot Pruning vs Iterative Pruning
As detailed in Section 3.4, Taco-SVD employs a layer-wise backward compression strategy to compress redundant layers. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. Table 5 shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22
Table 5: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
5 Discussion
5.1 How Taco-SVD Stabilizes Layer Removal and Enhances Model Consistency
Instabilities Caused by Layer Removal
Layer removal can disrupt the internal consistency of a model, leading to performance degradation. In response, the model attempts to compensate for the loss of removed layers, triggering a process we refer to as self-repair (Rushing and Nanda, 2024). While this self-repair mechanism helps the model maintain partial functionality, it is often noisy and incomplete. The process cannot fully recover the removed components, resulting in higher perplexity and reduced cross-task generalization. This phenomenon is task-dependent, with certain tasks being more sensitive to the loss of critical task-relevant transformations.
Taco-SVD as a Stabilizing Solution
Taco-SVD effectively addresses these issues by preserving key task-critical singular value directions, reducing the need for self-repair and stabilizing the model’s internal consistency. Specifically, Taco-SVD:
  • Retains task-critical singular values, minimizing disruption to the residual stream. This helps maintain the original structure and function of the model, reducing perplexity and improving performance across diverse tasks.
  • Mitigates the reliance on self-repair, ensuring that the model does not need to rely on incomplete compensations. By aligning the retained singular values with task-specific objectives, Taco-SVD facilitates smoother adjustments that preserve both task performance and internal consistency.
5.2 Taco-SVD vs LoRA vs LLM-Streamline
Limitations of fine-tuning Pruned LLMs with LoRA
Parameter-efficient fine-tuning methods such as LoRA introduce additional parameters across the remaining layers to compensate for the removed layers. However, training the pruned model with LoRA is a process of redistributing the function of the pruned layers (Chen et al., 2024), these updates often introduce intruder dimensions, which may lead to potential issues such as Spectral Dilution (Shuttleworth et al., 2024): LoRA’s weight updates alter the original singular value spectrum, the new singular vectors added by LoRA are often orthogonal to the original pre-trained singular vectors, which diminishes the contribution of task-agnostic features learned during pre-training.
Limitations of Layer Replacement with Dense Modules
Methods like LLM-Streamline (Chen et al., 2024) train a dense module from scratch to approximate the functionality of redundant layers within LLMs. However, dense modules possess high-dimensional internal structures, which may cause overfitting in scenarios with limited data.
Taco-SVD’s Generalization Advantage via Low-Rank Learning
In contrast, Taco-SVD avoids such issues by preserving the original spectral structure while aligning it with task-specific objectives.
  • Taco-SVD identifies and retains the singular value directions in the redundant layer that are most critical to task performance. It maintains the original pre-trained singular value spectrum, preventing the introduction of orthogonal updates to the original singular vectors in the remaining layers.
  • Previous literature has demonstrated the effectiveness and practicality of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022; Yaras et al., 2024). Therefore, instead of training a dense module from scratch, Taco-SVD fine-tunes only the minimal task-aware subspaces in redundant layers. It can avoid overfitting by constraining the model to learn within a low-rank structure.
6 Conclusion
We propose Taco-SVD, a task-aware framework for compressing large language models by retaining task-critical singular value directions. Taco-SVD mitigates the disruptions caused by traditional layer removal and avoids the need for computationally expensive fine-tuning compensation by preserving key task-critical transformations. Experimental results show that Taco-SVD achieves lower perplexity, better task performance, and improved cross-task robustness with minimal resource overhead. This work highlights the effectiveness of aligning low-rank decompositions with task-specific objectives for efficient and robust model compression.
7 Limitations
While Taco-SVD demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
Dependence on Calibration Data Taco-SVD relies on calibration data for singular value gradient attribution. In our experiments, using as few as 64 calibration samples from the general-purpose WikiText dataset achieved strong performance across multiple tasks. However, while this approach works well in these settings, we have not yet explored whether using domain-specific calibration data could yield better results in specialized domains. Further testing with such data may provide valuable insights into its potential advantages.
Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, Taco-SVD retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal. This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios. Further reduction techniques may complement Taco-SVD in such scenarios.
Scalability to Extremely Large Models While Taco-SVD performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges. Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
Ethical Considerations
Our research adheres to the ACL Code of Ethics, ensuring transparency, responsible use of data, and consideration of potential social impacts. All datasets used in this work are publicly available and have been appropriately cited, ensuring compliance with data usage agreements and privacy regulations.
While Taco-SVD is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight. To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
Moreover, we acknowledge the computational resources required for training and testing large language models. To minimize environmental impact, we conducted experiments on energy-efficient hardware (NVIDIA A100 GPUs) and report our computational cost transparently. Further details can be found in the Appendix.
Appendix A Appendix
A.1 The gradient of singular values
For a weight matrix $W\in\mathbb{R}^{m\times n}$ in the selected redundant layers, its differential form can be expressed as:
$\partial W=\partial U\Sigma V^{T}+U\partial\Sigma V^{T}+U\Sigma\partial V^{T}$
$U^{T}\partial WV=U^{T}\partial U\Sigma+\partial\Sigma+\Sigma V^{T}\partial V$
Since both $U$ and $V$ are orthogonal matrices, we have:
$U^{T}U=I_{m},\quad V^{T}V=I_{n}$
$\partial U^{T}U+U^{T}\partial U=O_{m},\quad\partial V^{T}V+V^{T}\partial V=O_{n}$
This implies that $U^{T}dU$ and $dV^{T}V$ are asymmetric matrices. Therefore, the diagonal elements of $U^{T}dU\Sigma$ and $\Sigma V^{T}dV$ are zero, leading to the diagonal elements of $U^{T}\partial WV$ being:
$I_{k}\odot U^{T}\partial WV=\partial\Sigma$
where $I_{k}$ represents the $k\times k$ identity matrix, $\odot$ denotes element-wise multiplication.
For a singular value $\sigma_{i}$, its differential form can be written as:
$\partial{\sigma_{i}}=u_{i}^{T}\partial Wv_{i}$
Since $\sigma_{i}$ is a scalar, we have:
$\partial\sigma_{i}$ $=\text{tr}(\partial\sigma_{i})$
$=\text{tr}(u_{i}^{T}\partial Wv_{i})$
$=\text{tr}[(u_{i}v_{i}^{T})^{T}\partial W]$
thereby, the derivative of $\sigma_{i}$ with respect to W is:
$\frac{\partial\sigma_{i}}{\partial W}=u_{i}v_{i}^{T}$
For a calibration dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
$\frac{\partial L}{\partial\sigma_{i}}=u_{i}^{T}\frac{\partial L}{\partial W}v_{i}$
Then, for all the singular values $\Sigma$, we have:
$\frac{\partial L}{\partial\Sigma}=I_{k}\odot U^{T}\left(\frac{\partial L}{\partial W}\right)V$
A.2 Layer-wise Low-Rank Decomposition
This section provides the implementation of Layer-wise Low-Rank Decomposition, as detailed in Algorithm 1.
Algorithm 1 Task-Aware SVD for LLM Compression
1:Model $M$ with $N$ layers, Calibration Dataset $D$, Retain Ratio $r$
2:$\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$ for weight matrices in self-attention and MLP blocks
3:Compressed Model $\tilde{M}$
4:Step 1: Redundant Layer Selection
5:for $i=1$ to $N$ do
6:     $\text{cos}_{i}\leftarrow\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$
7:end for
8:Select $L$ layers with highest cosine similarity.
9:Step 2: Task-aware SVD Compression:
10:for each layer $l\in L$, starting from the last layer do
11:     $\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$
12:     $\text{W}\leftarrow\mathbf{U}\Sigma\mathbf{V^{T}}$
13:     $\text{G}\leftarrow\frac{\partial\mathcal{L}_{\text{}}}{\partial W}$   # compute gradient matrices using $D$
14:     $\frac{\partial\mathcal{L_{\text{}}}}{\partial\sigma_{i}}\leftarrow\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$   # Compute gradients for singular values
15:     $\tilde{W}\leftarrow\tilde{U}\tilde{\Sigma}\tilde{V^{T}}$   # Truncate singular values based on Eq.5, retain only $r\%$ parameters per matrix
16:end for
17:return $\tilde{M}$
Figure 5: Validation loss curves of different methods on the Alpaca dataset (1 epoch)
A.3 Experimental Setup and Hyperparameters Configuration
To ensure a fair comparison, all experimental setups are consistent across all methods. Below we describe the experimental setup and hyperparameters configuration in detail.
Training-free Setting
We report the results of Taco-SVD and other training-free baselines under the compression ratio of 20% in Table 1. Specifically, we selected the same 256 samples from the WikiText-2 dataset as calibration data.
Training-based Setting
Table 6 provides the detailed configurations of training-based methods under the compression ratio of 20%. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled.
HyperParameters Setting
Dataset Alpaca
Huggingface Dataset Path yahma/alpaca-cleaned
Batch Size 32
Micro Batch Size 4
Epochs 1
Learning Rate 3.00E-04
Max Length 256
Train on Inputs TRUE
Add EOS Token FALSE
LoRA-Rank 256
LoRA-Alpha 16
LoRA-Dropout 0.05
LoRA-Target-Modules
q_proj, k_proj, v_proj, o_proj,
up_proj, down_proj, gate_proj
Prompt-Template Alpaca Template
Table 6: Experimental setup and hyperparameter configurations.
Method Ratio WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Dense 0% 5.47 20.83 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LaCo 18% 39.43 93.98 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
SliceGPT 21.50% 8.63 99.25 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
ASVD 20% 9.7 88.39 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
Taco-SVD 21.60% 16.12 44.09 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
LaCo 24% 50.33 102.84 25 45.33 60.06 38.3 29.69 65.51 23.99 41.13
SliceGPT 25.10% 9.54 118.17 20.6 50.91 59.63 35.66 26.41 61.62 23.43 39.75
ASVD 25% 53.48 374.45 17 39.35 53.12 32.64 22.78 61.48 22.95 35.62
Taco-SVD 25% 30.2 60.74 21.6 46.59 61.8 40.41 28.84 66.1 23.18 41.22
LaCo 30% 94.5 122.09 22 42.26 59.43 35.7 28.41 62.89 24.32 39.29
SliceGPT 29.10% 10.91 145.24 18.8 47.56 58.64 34.18 25.34 59.58 23.15 38.18
ASVD 30% 1877.63 3340.65 14.4 27.78 50.59 26.37 21.84 51.58 20.88 30.49
Taco-SVD 30% 35.38 88.96 19.4 42.72 61.17 38.43 27.73 63.98 23.18 39.52
LaCo 36% 138.82 164.02 19 38.34 58.96 33.71 27.13 60.55 23.55 37.32
SliceGPT 33.20% 12.8 176.54 17.8 44.23 57.61 32.28 23.55 58.59 22.68 36.68
ASVD 35% 2599.79 6529.72 13.6 26.81 50.83 25.94 23.12 53.59 19.33 30.46
Taco-SVD 35.00% 68.43 125.79 18.8 40.25 60.35 36.77 27.18 61.11 22.63 38.16
LaCo 39% 386.55 289.98 18.4 33.96 56.99 31.34 26.88 58.6 23.45 35.66
SliceGPT 41.40% 19.1 257.66 15.8 35.94 53.43 29.62 20.82 55.98 22.01 33.37
ASVD 40% NaN NaN 13.4 25.76 51.3 26.1 21.67 54.19 20.64 30.44
Taco-SVD 40.10% 80.42 247.55 18 36.83 59.59 33.31 26.62 59.63 21.88 36.55
Table 7: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and training-free baselines under 20% to 40% compression. Bold values indicate the best performance.
A.4 Validation Loss Curves
We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods. As shown in Figure 5, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
A.5 Results under High Compression Ratio
In this section, we present the results of Taco-SVD compared with the training-free baselines using LLaMA2-7B under high compression ratios, as shown in Table 7. The results demonstrate that our proposed method consistently outperforms the baselines, even at high compression ratios of up to 40%.
A.6 Robustness of Taco-SVD towards different Calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the number of calibration data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of Taco-SVD to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 6, we can observe that Taco-SVD consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of calibration data. Tables 8 and 9 summarize the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
Figure 6: Performance of Taco-SVD on LLaMA3.1-8B-Instruct under 20% compression using (a) different calibration datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. Taco-SVD demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
WikiText-2 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
C4 40.54 71.42 24 57.91 67.72 42.11 36.69 67.25 27.5 46.17
Table 8: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by Taco-SVD under 20% compression using 512 samples from WikiText-2 and C4 as calibration datasets.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
WikiText-2-64 46.5 86.51 22.6 59.97 69.3 44.36 37.29 68.5 27.37 47.06
WikiText-2-128 39.91 76.41 22.8 60.23 69.93 44.24 35.92 67.9 27.5 46.93
WikiText-2-256 38.73 79.13 21.8 59.89 70.24 44.23 36.26 67.19 27.07 46.67
WikiText-2-512 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 9: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by Taco-SVD under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.