### v2

# Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression

Abstract
Layer removal is an effective technique for compressing large language models (LLMs) by reducing redundancy and improving inference efficiency. However, indiscriminate pruning disrupts representation stability, leading to performance degradation. We propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), which preserves representation-critical singular values to mitigate these effects. Unlike direct layer removal, GRASP leverages gradient-based attribution on a syntax- and semantics-rich dataset to guide the selection of representation-critical singular values. By selectively applying singular value decomposition (SVD) to affected layers, GRASP achieves efficient compression while maintaining representation stability with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods in perplexity and downstream task performance. The source code is available at GRASP.
[…]
Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
Kainan Liu1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1Ping An Technology (Shenzhen) Co., Ltd., China 2The Hong Kong University of Science and Technology (Guangzhou) {zhangyong203, chengning211}@pingan.com.cn
footnotetext: † Equal contribution.footnotetext: * Corresponding author.footnotetext: This work was done during Kainan Liu’s internship at Ping An Technology.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023b). However, their massive parameter sizes pose computational and memory challenges, hindering deployment on resource-limited devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
Figure 1: Zero-shot performance of GRASP and pure layer removal under 20% compression ratio. Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
Despite its simplicity and efficiency, layer removal introduces two main challenges, as shown in Figure 1. First, while task accuracy may remain stable in some cases, the perplexity (PPL) often degrades significantly, reflecting disruptions in the model’s internal representation consistency. This degradation arises from impaired information flow and misaligned intermediate representations (Liang et al., 2024). The sharp increase in PPL suggests that layer removal eliminates critical transformations necessary for maintaining coherent representation structures, potentially limiting the model’s generalization on complex or unseen tasks. Second, the impact of layer removal varies significantly across models. For example, under the same pruning ratio, LLaMA2 (Touvron et al., 2023b) maintains reasonable performance, while LLaMA3 (Dubey et al., 2024) experiences a steep performance drop. This discrepancy suggests that layer removal affects models unevenly, as certain structural transformations are more essential for maintaining stable representations, leading to performance inconsistencies across models.
Our analysis suggests that these issues arise from the removal of transformations that play an essential role in preserving internal consistency and maintaining structured representations. In this paper, we introduce GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel hybrid pruning framework that integrates layer removal with gradient-based singular value selection for efficient LLM compression. Unlike conventional pruning techniques that either remove full layers or apply global SVD compression, GRASP selectively applies singular value decomposition (SVD) only to layers affected by pruning. A key component of GRASP is its use of a small, syntax- and semantics-rich dataset (e.g., WikiText-2), which serves as a representation evaluation dataset to guide the selection of singular values. Instead of relying solely on magnitude-based ranking, GRASP leverages gradient attribution computed from this dataset to dynamically identify and retain singular values critical for preserving language representations. This ensures that despite layer removal, the model retains essential syntactic and semantic structures, minimizing degradation in perplexity and downstream task performance.
We evaluate GRASP on 10 datasets and five models from two distinct LLM families (LLaMA, Mistral). Extensive experiments validate the effectiveness of GRASP. It achieves compression rates comparable to pure layer removal while significantly improving perplexity and task performance across a wide rage of downstream tasks. Notably, GRASP maintains strong robustness across different models without requiring additional training. Furthermore, it achieves strong performance with minimal calibration data—using only 64 samples from a general-purpose, unlabeled dataset, demonstrating the efficiency of our approach. Additionally, post-training compensation further enhances GRASP’s performance, highlighting its adaptability in compressing large language models while preserving their linguistic capabilities.
Figure 2: (A): Overview of GRASP. The figure illustrates how GRASP uses gradient attribution to identify representation-critical singular values and applies SVD to compress redundant layers while preserving representation-relevant transformations. (B): SVD-based LLM compression and (C): Layer removal compression methods
2 Related Work
2.1 Layer Removal-based LLM Compression
Layer removal is a structured pruning technique that eliminates redundant layers within Large Language Models (LLMs), aiming to reduce the model’s computational and memory requirements. Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers. SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs. LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
Although effective in reducing model size, layer removal disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al. (2024). To mitigate this, post-training compensation methods have been proposed. Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities. However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024). LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers. While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
2.2 SVD-based LLM Compression
Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss. Recent methods have enhanced SVD to reduce LLM compression error. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a representation evaluation dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
GRASP differs from these methods in three key aspects. First, rather than minimizing F2 norm reconstruction loss, GRASP optimizes a representation learning attribution loss, computed on a small representation evaluation dataset to ensure selection aligns with model representation rather than weight approximation. Second, instead of using a top-k singular value selection strategy, GRASP leverages gradient attribution to dynamically retain representation-critical singular values, which are not necessarily the largest in magnitude. Third, while SVD-based pruning is typically applied to all layers, GRASP is selectively used only in pruned layers, ensuring that layer removal does not disrupt critical transformations.
3 Method
3.1 Overview and Motivation
In this paper, we propose GRASP, a novel hybrid gradient-guided compression framework designed to compress redundant layers in LLMs while preserving representation-critical transformations. Figure 2 provides an overview of GRASP. The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) selecting and retaining representation-critical singular values and their corresponding singular vectors using gradient-based attribution. Below, we describe each step in detail.
3.2 Redundant Layer Selection
The first step in GRASP is identifying redundant layers. These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance. Following prior works (Song et al., 2024; Chen et al., 2024), we use cosine similarity to quantify the degree of transformation in a given layer.
For a transformer layer with input hidden state $H_{i}\in\mathbb{R}^{d}$ and output hidden state $H_{i+1}\in\mathbb{R}^{d}$, the cosine similarity is computed as:
$\text{cos}(H_{i},H_{i+1})=\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$ (1)
A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant. Instead of directly removing these layers, GRASP compresses weight matrices with a gradient-based approach to retain critical internal transformations.
3.3 Gradient-Guided Singular Value Selection
3.3.1 Representation Learning Attribution Loss (RLAL)
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone. Unlike traditional SVD-based methods that select top-$k$ singular values using a reconstruction loss measured by the Frobenius norm ($F_{2}$ norm), GRASP utilizes a Representation Learning Attribution Loss (RLAL), designed specifically for general representation capability modeling.
RLAL is computed on a small, unlabeled representation evaluation dataset derived from corpora rich in syntactic and semantic structures, such as WikiText-2 and similar datasets. These datasets ensure that the retained singular values are those most crucial for maintaining the model’s ability to capture fundamental language structures, including syntax and semantics, even after compression. Unlike task-specific objectives, RLAL focuses on preserving linguistic consistency and general-purpose representation learning rather than optimizing for any particular downstream application.
Mathematically, RLAL is formulated as:
$\mathcal{L}_{\text{RLAL}}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (2)
where $y_{t}$ is the target token at time step $t$, $x_{\leq t}$ denotes the sequence of preceding tokens, and $P(y_{t}\mid x_{\leq t})$ represents the predicted probability of the next token given the context. This loss function aligns with the standard language modeling loss, ensuring that the model learns meaningful representations.
3.3.2 Singular Value Decomposition (SVD)
For a redundant layer’s weight matrix $W_{\text{redundant}}\in\mathbb{R}^{m\times n}$, SVD decomposes $W_{\text{redundant}}$ as:
$W_{\text{redundant}}=U\Sigma V^{T}$ (3)
where $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{k})$ contains the singular values ($\sigma_{i}\geq 0$), and $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, respectively.
3.3.3 Representation-Critical Singular Value Importance
To assess the representation-critical importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Representation-Critical Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the $\mathcal{L}_{\text{RLAL}}$. Using its first-order Taylor expansion, the alteration in the $\mathcal{L}_{\text{RLAL}}$ caused by $\sigma_{i}$ is approximated as:
$\Delta\mathcal{L}_{\text{RLAL}}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (4)
where $\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the representation learning attribution loss , and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$ onto the singular directions:
$\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (5)
where $G=\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial W}$ is the gradient of the $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$. The detailed proof can be found in the Appendix A.1.
Integrating the gradient sensitivity with the magnitude of $\sigma_{i}$, we define $\Phi(\sigma_{i})$ as:
$\Phi(\sigma_{i})=\sigma_{i}\cdot\left(\mathbf{u}_{i}^{T}G\mathbf{v}_{i}\right)$ (6)
where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to representation performance, considering both its magnitude and alignment with the gradient of the general-purpose model loss. A larger value of $\Phi(\sigma_{i})$ indicates greater importance.
We rank singular values in descending order of representation relevance based on $\Phi(\sigma_{i})$. Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
$W_{\text{RLAL}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (7)
This ensures that $W_{\text{RLAL}}$ preserves the transformations most critical to representation stability, achieving efficient compression while maintaining representation consistency.
3.4 Layer-wise Low-Rank Decomposition
GRASP applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward. For each redundant layer, weight matrices are compressed using Gradient-Guided Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$. This ensures a low-rank approximation that preserves representation-critical transformations.
By processing one layer at a time, this approach significantly reduces memory overhead, avoiding the need to load the entire model into memory. Additionally, the backward order ensures that compressing deeper layers does not disrupt the consistency of earlier layers, preserving the overall stability of the model. The detailed implementation for Layer-wise Low-Rank Decomposition is provided in Algorithm 1 in the Appendix A.2.
4 Experiments
In this section, we conduct comprehensive experiments to evaluate GRASP from four key perspectives. (1) We first compare our method with several popular pruning-based LLM compression methods to demonstrate its effectiveness (Section 4.2). (2) Next, we compare GRASP with SVD-based LLM compression methods under different compression ratios with different LLMs (Section 4.3). (3) Then, we further evaluate GRASP on the LongBench (Bai et al., 2023) datasets to assess its performance on long-form text processing tasks (Section 4.4). (4) Finally, we analyze the inference speed-up achieved by GRASP (Section 4.5).
4.1 Experimental Setup
Below we detail the models, benchmarks and implementation details, with more experimental setups provided in Appendix A.3 due to page limit.
Models
We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023b), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral family models Mistral-7B (Jiang et al., 2023).
Benchmark
10 datasets are used as evaluation benchmarks including two pre-training datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), seven common sense reasoning datasets include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018), MathQA (Amini et al., 2019) and one generation benchmark GSM8K (Cobbe et al., 2021). Additionally, we include LongBench (Bai et al., 2023) to assess LLM’s performance on long-form text processing after pruning.
Implementation Details
For all baselines, we randomly sample 512 instances from the WikiText-2 dataset as the representation evaluation dataset. For pruning-based with post-training compensation baselines, all models are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. Further experimental details can be found in Appendix A.3.
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
ShortGPT 23.95 80.70 0.46 18.45 57.58 0.43 27.26 59.14 0.44 140.47 262.47 0.44
LaCo 244.53 524.26 0.34 39.43 93.98 0.43 38.27 125.98 0.42 753.98 1577.44 0.43
SliceGPT 9.04 107.65 0.39 8.63 99.25 0.41 7.44 99.77 0.45 52.78 460.73 0.35
Ours 18.42 77.98 0.46 16.12 44.09 0.44 11.1 56.57 0.47 37.86 63.97 0.47
Table 1: Perplexity($\downarrow$) of GRASP and pruning-based without post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio. The best performance is marked in bold.
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
LoRA 11.56 34.05 0.51 11.7 35.39 0.46 8.46 58.15 0.50 14.56 21.24 0.52
LS-FFN 16.12 59.98 0.49 13.94 37.43 0.46 8.11 76.3 0.50 16.79 24.76 0.51
LS-Layer 14.67 54.31 0.49 11.57 32.62 0.47 8.23 77.04 0.50 14.27 20.85 0.53
Ours* 11.62 44.14 0.51 9.59 30.38 0.48 7.52 38.31 0.51 14.13 20.99 0.53
Table 2: Perplexity($\downarrow$) of GRASP (Ours* denotes the version with post-training compensation) and pruning-based with post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. The best performance is marked in bold.
Ratio Method WikiText-2 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA GSM8K Average Percentage
0% Original 5.68 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.09 0.46 100.0
20% FWSVD 1727 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.00 0.28 60.8
ASVD 11.14 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.04 0.38 82.6
SVD-LLM 7.94 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.05 0.39 84.7
Ours 14.79 0.22 0.52 0.64 0.43 0.32 0.70 0.24 0.03 0.39 84.7
Ours* 10.19 0.24 0.59 0.63 0.5 0.35 0.73 0.25 0.04 0.42 91.3
30% FWSVD 20127 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.00 0.26 56.5
ASVD 51 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.00 0.33 71.7
SVD-LLM 9.56 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.03 0.35 76.1
Ours 30.48 0.19 0.42 0.62 0.39 0.28 0.64 0.23 0.02 0.35 76.1
Ours* 12.63 0.24 0.54 0.64 0.46 0.32 0.69 0.24 0.04 0.40 87.0
40% FWSVD 18156 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.00 0.27 58.7
ASVD 15358 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.00 0.26 56.5
SVD-LLM 13.11 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.02 0.33 71.7
Ours 62.34 0.18 0.37 0.57 0.35 0.27 0.61 0.21 0.01 0.32 69.6
Ours* 16.25 0.22 0.49 0.63 0.43 0.3 0.68 0.23 0.02 0.38 82.6
50% FWSVD 24391 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.00 0.26 56.5
ASVD 15358 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.00 0.26 56.5
SVD-LLM 23.97 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.00 0.29 63.0
Ours 102.58 0.13 0.29 0.53 0.28 0.23 0.53 0.20 0.01 0.28 60.9
Ours* 23.69 0.18 0.4 0.56 0.35 0.26 0.61 0.21 0.02 0.32 69.6
[…]
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
Original 5.25 0.59 5.68 0.52 5.47 0.53 7.21 0.60
FWSVD 6357 0.32 1727 0.32 2360 0.31 3256.7 0.29
ASVD 19.28 0.4 11.14 0.44 9.70 0.46 2443.99 0.30
SVD-LLM 10.21 0.41 7.94 0.44 8.50 0.47 - -
Ours 18.42 0.45 14.79 0.44 16.12 0.44 37.86 0.47
Ours* 11.62 0.51 10.19 0.47 9.59 0.48 14.13 0.53
Table 4: Perplexity($\downarrow$) of GRASP and SVD-based baselines on the WikiText-2 datasets and the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio. "†" indicates that we refer to the results in the original paper. The best performance is marked in bold.
Model
Single-Doc
QA
Multi-Doc
QA
Summarization
Task
Few-shot
Learning
Synthetic
Task
Code
Completion
Average
LLaMA3.1-8B-Instruct 12.26 16.16 22.19 42.52 9.49 53.91 26.09
GRASP* 21.69 24.41 20.03 45.84 9.67 35.92 26.26
Table 5: Evaluations of LLaMA3.1-8B-Instruct and its compressed version of GRASP* under 20% compression ratio on LongBench benchmarks. The average results across six task categories are shown in this table, with additional details provided in the Appendix A.7.
4.2 Comparison with Pruning-based LLM Compression Methods
Post-training compensation plays a crucial role in restoring the performance of compressed LLMs. To ensure a fair comparison, we compare GRASP with two categories of pruning-based compression methods. (1) Pruning without post-training compensation, including two layer removal methods, ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024), and state-of-the-art structured pruning method: SliceGPT (Ashkboos et al., 2024). (2) Pruning with post-training compensation: ShortGPT with LoRA (Hu et al., 2022) and LLM-Streamline (Chen et al., 2024). We provide a detailed comparison of these pruning-based LLM compression methods in Appendix A.4.
Pruning without post-training compensation
We present the perplexity on WikiText-2 and PTB datasets, as well as the average accuracy on common sense reasoning benchmarks for different methods in Table 1. The results demonstrate that GRASP consistently achieves strong performance across all models and tasks. Specifically, in common sense reasoning tasks, GRASP surpasses SliceGPT by 15% in accuracy on Mistral-7B and by 34% on LLaMA3.1-8B-Instruct. For language generation tasks, while SliceGPT achieves lower perplexity on the WikiText-2 dataset, GRASP consistently outperforms the baselines on the PTB dataset. We hypothesize this discrepancy arises from the potential overfitting problem of baselines to the calibration dataset. Notably, GRASP exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse models. Detailed results for each common sense reasoning benchmark are provided in Appendix A.5.
Pruning with Post-Training Compensation
We further demonstrate that GRASP provides a better initialization for post-training compensation, enabling more accurate recovery of the compressed model compared to other methods. To ensure a fair comparison, we constrain the number of trainable parameters to remain approximately the same across all methods by retaining only 10% of the parameters in each redundant layer and allowing only these parameters to be trainable. As shown in Table 2, GRASP consistently outperforms LLM-Streamline across different LLMs. Furthermore, we present the validation loss curves for each method during post-training in Appendix A.6, which show that GRASP GRASP enables more stable and faster convergence. More Detailed results are provided in Appendix A.5.
4.3 Comparison with SVD-based LLM Compression Methods
In this section, we compare the performance of GRASP with state-of-the-art SVD-based LLM compression methods—FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024)—from two distinct perspectives: (1) performance under different compression ratios and (2) performance across different LLMs.
Performance under Different Compression Ratios.
To evaluate the performance and stability of our proposed method under various compression ratios, we compare GRASP with the SVD-based baselines on LLaMA-7B under compression ratios ranging from 20% to 50%. Table 3 summarizes the results for different methods. The results demonstrate that our proposed GRASP consistently outperforms the baseline methods on most benchmarks, except for a slightly higher perplexity on WikiText-2 compared to SVD-LLM. Specifically, GRASP retains more than 91% of the original performance at a 20% compression ratio and 87% under the compression ratio of 30%. More importantly, with fast and resource-efficient post-training, GRASP enables rapid accuracy recovery, achieving 70% of the original model performance even at a 50% compression ratio.
Performance on different LLMs.
To evaluate the generalizability of GRASP across different LLMs, we compare its performance with the SVD-based baselines under 20% compression ratio on four models from two distinct LLM families-LLaMA 2-7B, LLaMA 2-13B, LLaMA 3.1-8B-Instruct and Mistral-7B. The results, presented in Table 4, indicate that while existing methods achieve competitive performance on the LLaMA 2 series, they exhibit a sharp performance drop on other architectures, particularly on LLaMA 3. In contrast, GRASP consistently outperforms baseline methods across all architectures and demonstrates greater stability across different LLMs.
4.4 Performance in Long-Form Text Processing
In this section, we evaluate the performance of GRASP on LongBench (Bai et al., 2023) to assess its performance on long-form reasoning and complex generative tasks. The average performance across six task categories: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion are summarized in Table 5, with detailed results provided in the Appendix A.7. As shown, the compressed version by GRASP under 20% compression ratio still maintains strong performance and even surpasses the original LLaMA3.1-8B-Instruct model in single-document QA, multi-document QA, and few-shot learning tasks. This demonstrates the effectiveness of GRASP in rapidly recovering accuracy and generative capability through post-training compensation after pruning.
4.5 Compression Costs and Inference Efficiency of GRASP
GRASP can compress LLMs at very low cost while improve the inference efficiency on real hardware. To evaluate its acceleration effect, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by GRASP with different sequence lengths and batch sizes. As shown in Figure 3, GRASP consistently consistently improves generation speed and achieves an acceleration effect comparable to pure layer removal. It is important to note that although GRASP retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance. Additionally, we also measure the compression time and peak memory consumption of GRASP when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. The detailed results are provided in Appendix A.8.
Figure 3: Throughput of LLaMA2-7B and GRASP compressed model under 25% compression ratio on a single A100 GPU. Top: Throughput across different sequence lengths (batch size = 32). Bottom: Throughput across different batch sizes (sequence length = 32).
4.6 Ablation Study
Calibration Data
We conducted ablation studies to investigate the impact of representation evaluation dataset and the number of data used for singular value gradient attribution. Table 6 summarizes GRASP’s performance when compressing LLaMA3.1-8B-Instruct using different representation evaluation datasets and varying numbers of data from WikiText-2. The results demonstrate that GRASP is robust to the choice of both the representation evaluation dataset and the data number. Notably, GRASP achieves competitive performance with just 64 samples, highlighting its ability to perform effective compression with minimal data. Detailed results can be found in Appendix A.9.
Ablation Type
Calibration
Dataset
WikiText-2 PTB
Average
Accuracy
Varying Dataset WikiText-2 37.86 63.97 47.12
C4 40.54 71.42 46.17
Varying Number 64 46.5 86.51 47.06
128 39.91 76.41 46.93
256 38.73 79.13 46.67
512 37.86 63.97 47.12
Table 6: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
One-shot Pruning vs Iterative Pruning
As detailed in Section 3.4, GRASP employs a layer-wise backward compression strategy to compress redundant layers. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. We present the ablation results in Table 12 in Appendix A.10, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
5 Discussion
GRASP preserves representation stability.
Layer removal disrupts the internal spectral structure of LLMs, often introducing inconsistencies in the learned representations. These disruptions lead to degradation in perplexity and loss of linguistic coherence, as the model struggles to compensate for missing transformations. GRASP stabilizes the compression process by selectively retaining representation-critical singular values, ensuring that the model’s residual stream remains structurally consistent. Unlike direct layer removal, which forces the model to undergo self-repair (Rushing and Nanda, 2024), GRASP proactively mitigates this instability by preserving the most influential spectral components, thereby reducing the burden on later layers to reconstruct lost information.
GRASP accelerates adaptation via low-rank learning.
Fine-tuning pruned models using parameter-efficient methods such as LoRA often introduces intruder dimensions into the pre-trained singular spectrum (Shuttleworth et al., 2024), leading to spectral dilution that weakens the contribution of pre-trained features. Similarly, replacing layers with dense modules, as in LLM-Streamline (Chen et al., 2024), requires training high-dimensional structures from scratch, which increases the risk of overfitting when data is limited. In contrast, GRASP identifies and fine-tunes the most essential low-rank subspaces within redundant layers, preserving the original spectral alignment while adapting efficiently to new tasks. Prior research has established the effectiveness of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022), and GRASP leverages this principle to optimize task-specific performance while maintaining pre-trained feature integrity.
6 Conclusion
We present GRASP, a novel framework for compressing large language models by selectively retaining representation-critical singular value directions. By mitigating the disruptions caused by layer removal, GRASP preserves essential transformations necessary for maintaining model stability and generalization. Experimental results demonstrate that GRASP achieves efficient and robust compression, improving perplexity and downstream task performance compared to existing methods while maintaining minimal computational overhead.
7 Limitations
While GRASP demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
Dependence on Calibration Data GRASP relies on representation evaluation dataset for singular value gradient attribution. In our experiments, using as few as 64 samples from the WikiText dataset achieved strong performance across multiple tasks. However, while this approach works well in these settings, we have not yet explored whether using domain-specific dataset could yield better results in specialized domains. Further testing with such data may provide valuable insights into its potential advantages.
Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, GRASP retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal. This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios. Further reduction techniques may complement GRASP in such scenarios.
Scalability to Extremely Large Models While GRASP performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges. Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
Ethical Considerations
Our research adheres to the ACL Code of Ethics, ensuring transparency, responsible use of data, and consideration of potential social impacts. All datasets used in this work are publicly available and have been appropriately cited, ensuring compliance with data usage agreements and privacy regulations.
While GRASP is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight. To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
Moreover, we acknowledge the computational resources required for training and testing large language models. To minimize environmental impact, we conducted experiments on energy-efficient hardware (NVIDIA A100 GPUs) and report our computational cost transparently. Further details can be found in the Appendix.
Appendix A Appendix
A.1 The gradient of singular values
For a weight matrix $W\in\mathbb{R}^{m\times n}$ in the selected redundant layers, its differential form can be expressed as:
$\partial W=\partial U\Sigma V^{T}+U\partial\Sigma V^{T}+U\Sigma\partial V^{T}$
$U^{T}\partial WV=U^{T}\partial U\Sigma+\partial\Sigma+\Sigma V^{T}\partial V$
Since both $U$ and $V$ are orthogonal matrices, we have:
$U^{T}U=I_{m},\quad V^{T}V=I_{n}$
$\partial U^{T}U+U^{T}\partial U=O_{m},\quad\partial V^{T}V+V^{T}\partial V=O_{n}$
This implies that $U^{T}dU$ and $dV^{T}V$ are asymmetric matrices. Therefore, the diagonal elements of $U^{T}dU\Sigma$ and $\Sigma V^{T}dV$ are zero, leading to the diagonal elements of $U^{T}\partial WV$ being:
$I_{k}\odot U^{T}\partial WV=\partial\Sigma$
where $I_{k}$ represents the $k\times k$ identity matrix, $\odot$ denotes element-wise multiplication.
For a singular value $\sigma_{i}$, its differential form can be written as:
$\partial{\sigma_{i}}=u_{i}^{T}\partial Wv_{i}$
Since $\sigma_{i}$ is a scalar, we have:
$\partial\sigma_{i}$ $=\text{tr}(\partial\sigma_{i})$
$=\text{tr}(u_{i}^{T}\partial Wv_{i})$
$=\text{tr}[(u_{i}v_{i}^{T})^{T}\partial W]$
thereby, the derivative of $\sigma_{i}$ with respect to W is:
$\frac{\partial\sigma_{i}}{\partial W}=u_{i}v_{i}^{T}$
For a representation evaluation dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
$\frac{\partial L}{\partial\sigma_{i}}=u_{i}^{T}\frac{\partial L}{\partial W}v_{i}$
Then, for all the singular values $\Sigma$, we have:
$\frac{\partial L}{\partial\Sigma}=I_{k}\odot U^{T}\left(\frac{\partial L}{\partial W}\right)V$
A.2 Layer-wise Low-Rank Decomposition
This section provides the implementation of Layer-wise Low-Rank Decomposition, as detailed in Algorithm 1.
Algorithm 1 Gradient-Guided SVD for LLM Compression
1:Model $M$ with $N$ layers, Representation Evaluation Dataset $D$, Retain Ratio $r$
2:$\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$ for weight matrices in self-attention and MLP blocks
3:Compressed Model $\tilde{M}$
4:Step 1: Redundant Layer Selection
5:for $i=1$ to $N$ do
6:     $\text{cos}_{i}\leftarrow\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$
7:end for
8:Select $L$ layers with highest cosine similarity.
9:Step 2: Gradient-guided SVD Compression:
10:for each layer $l\in L$, starting from the last layer do
11:     $\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$
12:     $\text{W}\leftarrow\mathbf{U}\Sigma\mathbf{V^{T}}$
13:     $\text{G}\leftarrow\frac{\partial\mathcal{L}_{\text{}}}{\partial W}$   # compute gradient matrices using $D$
14:     $\frac{\partial\mathcal{L_{\text{}}}}{\partial\sigma_{i}}\leftarrow\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$   # Compute gradients for singular values
15:     $\tilde{W}\leftarrow\tilde{U}\tilde{\Sigma}\tilde{V^{T}}$   # Truncate singular values based on Eq.6, retain only $r\%$ parameters per matrix
16:end for
17:return $\tilde{M}$
Method Metric Calibration Data Need Post-Training Training Data Training Dataset Size Training Module
ShortGPT Cosine Similarity WikiText-2 No None None None
LaCo Cosine Similarity WikiText-2 Optional Unpublished 1B Full Parameters
SliceGPT PCA
WikiText-2
Alpaca
Optional Alpaca 5k Full Parameters
Shortened LLaMA
Taylor
Perplexity
BookCorpus Optional
SlimPajama
Alpaca
627B
50k
Full Parameters
LoRA-Adapter
LLM-Streamline Cosine Similarity WikiText-2 Yes SlimPajama 30k
Lightweight
Network
GRASP Cosine Similarity WikiText-2 Optional Alpaca 50k
Low-rank
Modules
Table 7: Comparison of pruning-based LLM compression methods, where the metric indicates the criterion used to identify redundant modules. "Optional" refers to methods that can either work without post-training or recover performance through post-training. Shortened LLaMA consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
A.3 Experimental Setup and Hyperparameters Configuration
To ensure a fair comparison, all experimental setups are consistent across all methods. In the following, we describe the experimental setup and hyperparameters configuration in detail.
Metric
For the WikiText-2 and PTB datasets, we use perplexity ($\downarrow$) to assess model’s generative capacity. For the seven common sense reasoning datasets, performance is measured by accuracy ($\uparrow$), and for the GSM8K dataset, performance is evaluated using Exact Match Accuracy ($\uparrow$). All common sense reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024), while GSM8K is tested in the five-shot setting.
Hyperparameters Configurations
For post-training compensation, all models compressed by different methods are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$. All our experiments are conducted on a single A100 GPU with mixed precision enabled. Table 8 provides the detailed configurations of post-training compensation.
HyperParameters Setting
Dataset Alpaca
Huggingface Dataset Path yahma/alpaca-cleaned
Batch Size 32
Micro Batch Size 4
Epochs 1
Learning Rate 3.00E-04
Max Length 256
Train on Inputs TRUE
Add EOS Token FALSE
LoRA-Rank 256
LoRA-Alpha 16
LoRA-Dropout 0.05
LoRA-Target-Modules
q_proj, k_proj, v_proj, o_proj,
up_proj, down_proj, gate_proj
Prompt-Template Alpaca Template
Table 8: Experimental setup and hyperparameters configurations.
A.4 Comparison of Concurrent Pruning-based Methods
We provide a detailed comparison of concurrent pruning-based LLM compression methods, and the results are summarized in Table 7.
A.5 Detailed Results on Common Sense Reasoning Benchmarks
In this section, we provide the detailed results for each benchmark. Table 9 summarizes the results of pruning-based methods without post-training compensation on the common sense reasoning benchmark, while Table 10 presents the results of methods with post-training compensation.
Figure 4: Validation loss curves of different methods on the Alpaca dataset (1 epoch)
Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B-v0.1 Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
LaCo 21.1% 0.20 0.35 0.58 0.26 0.25 0.53 0.24 0.34
ShortGPT 21.1% 0.19 0.57 0.68 0.46 0.37 0.71 0.26 0.46
SliceGPT 20.0% 0.19 0.51 0.59 0.35 0.25 0.61 0.23 0.39
GRASP 20.0% 0.21 0.56 0.68 0.43 0.38 0.67 0.26 0.46
LLaMA2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
LaCo 18.1% 0.26 0.48 0.59 0.42 0.32 0.69 0.24 0.43
ShortGPT 21.1% 0.23 0.49 0.63 0.42 0.31 0.68 0.23 0.43
SliceGPT 21.5% 0.22 0.54 0.61 0.37 0.28 0.63 0.23 0.41
GRASP 21.6% 0.24 0.54 0.63 0.43 0.33 0.71 0.23 0.44
LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
LaCo 19.5% 0.28 0.52 0.63 0.43 0.33 0.70 0.25 0.45
ShortGPT 22.1% 0.24 0.50 0.63 0.46 0.33 0.7 0.24 0.44
SliceGPT 20.0% 0.29 0.59 0.65 0.39 0.32 0.64 0.24 0.45
GRASP 20.0% 0.26 0.61 0.66 0.47 0.35 0.73 0.24 0.47
LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
LaCo 19.0% 0.26 0.49 0.65 0.33 0.30 0.65 0.30 0.42
ShortGPT 21.7% 0.21 0.57 0.66 0.42 0.32 0.67 0.26 0.44
SliceGPT 20.0% 0.15 0.43 0.51 0.30 0.23 0.58 0.22 0.35
GRASP 20.0% 0.22 0.60 0.70 0.44 0.37 0.69 0.28 0.47
Table 9: Zero-shot performance of GRASP and pruning-based without post-training baselines under 20% compression ratio. Results are reported on seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
LoRA 25.1% 0.25 0.69 0.71 0.51 0.40 0.72 0.26 0.51
LS-FFN 25.2% 0.22 0.67 0.68 0.50 0.38 0.70 0.27 0.49
LS-Layer 24.1% 0.22 0.69 0.70 0.45 0.39 0.72 0.27 0.49
GRASP 24.3% 0.26 0.69 0.70 0.52 0.41 0.73 0.27 0.51
LLaMA-2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
LoRA 25.0% 0.26 0.58 0.63 0.48 0.33 0.71 0.24 0.46
LS-FFN 25.7% 0.25 0.56 0.63 0.48 0.33 0.71 0.25 0.46
LS-Layer 24.0% 0.25 0.60 0.62 0.50 0.35 0.73 0.24 0.47
GRASP 24.3% 0.28 0.61 0.65 0.50 0.36 0.72 0.25 0.48
LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
LoRA 25.0% 0.27 0.63 0.67 0.53 0.38 0.74 0.25 0.50
LS-FFN 25.4% 0.28 0.61 0.67 0.54 0.38 0.73 0.25 0.50
LS-Layer 24.4% 0.29 0.62 0.67 0.54 0.40 0.74 0.26 0.50
GRASP 24.2% 0.29 0.63 0.68 0.55 0.39 0.74 0.27 0.51
LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
LoRA 22.3% 0.28 0.70 0.69 0.51 0.41 0.73 0.30 0.52
LS-FFN 22.0% 0.26 0.70 0.70 0.49 0.41 0.72 0.30 0.51
LS-Layer 21.7% 0.27 0.74 0.69 0.51 0.43 0.73 0.32 0.53
GRASP 21.0% 0.28 0.72 0.70 0.52 0.43 0.74 0.31 0.53
Table 10: Performance comparison of GRASP and pruning-based with post-training baselines under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. Results are reported on seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
A.6 Validation Loss Curves
We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods. As shown in Figure 4, we can observe that GRASP has better initialization, leading to more stable and faster convergence during training.
A.7 Evaluation Results on LongBench
In this section, we present the detailed results of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench, which are presented in Table 13. The results illustrate that GRASP with post-training compensation still maintains superior performance on long-form reasoning and complex generative tasks.
A.8 Compression Costs of GRASP
To assess the compressed costs of GRASP, we measure its compression time and peak memory consumption when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. These metrics, summarized in Table 11, indicate that GRASP achieves a good balance between compression time and memory consumption.
Model
Average
Compression Time(h)
Compression Memory
Consumption(MiB)
LaCO 0.05 45046
ASVD 5.5 17632
SliceGPT 0.6 18195
GRASP 0.16 29050
Table 11: Compression time and peak compression memory usage of GRASP and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
A.9 Robustness of GRASP towards different Representation Evaluation Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of representation evaluation datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 5, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the representation evaluation dataset and the number of data. Tables 14 and 15 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different representation evaluation datasets (WikiText-2, C4) and varying numbers of calibration data.
A.10 Robustness of GRASP towards Different Pruning Strategy
In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 12
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22
Table 12: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
Model Summarization Few-shot Learning Synthetic Task
1-1 1-2 1-3 1-4 1-1 2-1 2-2 2-3 2-4 3-1 3-2
LLaMA3.1-8B-Instruct 28.35 20.04 25.85 14.51 63 51.3 39.29 16.5 1.64 9.67 17.17
GRASP 25.76 19.41 25.97 8.99 59.5 67.44 38.41 18 1 10 18
Model One-Doc QA Multi-Doc QA Code Completion Average
4-1 4-2 4-3 4-4 5-1 5-2 5-3 5-4 6-1 6-2 ALL
LLaMA3.1-8B-Instruct 9.96 6.04 18.75 14.27 11.27 11.57 7.65 34.16 56.47 51.34 26.09
GRASP 15.32 20.47 34.52 16.46 31.33 26.18 10.28 29.86 33.74 38.09 26.26
Table 13: Performance comparison of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench. The datasets are grouped as follows: (1-1 to 1-4) denote GovReport, QMSum, MultiNews, and VCSUM; (2-1 to 2-4) denote TREC, TriviaQA, SAMSum, and LSHT; (3-1 to 3-3) denote PassageCount, PassageRetrieval-en, and PassageRetrieval-zh; (4-1 to 4-4) denote NarrativeQA, Qasper, MultiFieldQA-en, and MultiFieldQA-zh; (5-1 to 5-4) denote HotpotQA, 2WikiMultihopQA, MuSiQue, and DuReader; (6-1 to 6-2) denote LCC and RepoBench-P.
Figure 5: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different representation evaluation datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
WikiText-2 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
C4 40.54 71.42 24 57.91 67.72 42.11 36.69 67.25 27.5 46.17
Table 14: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as representation evaluation datasets.
Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
WikiText-2-64 46.5 86.51 22.6 59.97 69.3 44.36 37.29 68.5 27.37 47.06
WikiText-2-128 39.91 76.41 22.8 60.23 69.93 44.24 35.92 67.9 27.5 46.93
WikiText-2-256 38.73 79.13 21.8 59.89 70.24 44.23 36.26 67.19 27.07 46.67
WikiText-2-512 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 15: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.