### v3

# GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression

Abstract
Recent studies have demonstrated that many layers are functionally redundant in large language models (LLMs), enabling model compression by removing these layers to reduce inference cost. While such approaches can improve efficiency, indiscriminate layer pruning often results in significant performance degradation. In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that mitigates this issue by preserving sensitivity-aware singular values. Unlike direct layer pruning, GRASP leverages gradient-based attribution on a small calibration dataset to adaptively identify and retain critical singular components. By replacing redundant layers with only a minimal set of parameters, GRASP achieves efficient compression while maintaining strong performance with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods, achieving 90% of the original model’s performance under 20% compression ratio. The source code is available at GRASP.
[…]
GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression
Kainan Liu1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1Ping An Technology (Shenzhen) Co., Ltd., China 2The Hong Kong University of Science and Technology (Guangzhou) {zhangyong203, chengning211}@pingan.com.cn
footnotetext: † Equal contribution.footnotetext: * Corresponding author.footnotetext: This work was done during Kainan Liu’s internship at Ping An Technology.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023b). However, their massive parameter sizes pose computational and memory challenges, hindering deployment on resource-limited devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods remove entire components such as neurons or layers to streamline the model, thereby achieving hardware efficiency and inference speedup.
Figure 1: Unlike conventional layer pruning, which either skips redundant layers—often causing moderate performance drops—or replaces them with lightweight modules that require additional training, GRASP (right) retains only the most critical 10% of parameters within the redundant layers, effectively preserving accuracy with minimal overhead.
In this work, we focus on structured layer pruning, which builds on prior findings that certain consecutive layers in large language models (LLMs) are functionally redundant. These findings have inspired approaches that either remove such layers entirely (Men et al., 2024, Yang et al., 2024, Kim et al., 2024) or replace them with lightweight modules (Chen et al., 2024) to reduce inference cost. Although layer removal is simple and computationally efficient, it often results in significant performance degradation. This degradation arises from disrupted information flow and misaligned intermediate representations (Liang et al., 2024), suggesting that layer removal eliminates certain important components that contribute meaningfully to maintaining model performance. Replacing redundant layers with lightweight modules can mitigate this issue to some extent, but such modules are typically randomly initialized, requiring substantial computational resources for training.
In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that replaces redundant layers with adaptive singular parameters for efficient LLM compression. Unlike direct layer removal, GRASP exploits the low-rank structure inherent in redundant layers, replacing the redundant layers with only a small subset of parameters while maintaining strong model performance. Specifically, GRASP operates in two key steps: First, it identifies layers suitable for pruning based on the cosine similarity of output hidden states between adjacent layers. Then, instead of relying on magnitude-based heuristics, GRASP leverages gradient attribution derived from a small calibration dataset to adaptively identify and retain the singular values most critical for downstream task performance.
To evaluate the effectiveness of GRASP, We conduct extensive experiments on 19 datasets and 5 models from two distinct LLM families (LLaMA and Mistral). Notably, GRASP achieves strong performance in a training-free setting, requiring no additional optimization. Furthermore, when post-training compensation is applied, only a small number of samples are needed to rapidly restore model performance. This efficiency arises from retaining critical components within redundant layers, rather than relying on randomly initialized replacements. Overall, this paper makes the following contributions:
  • We propose GRASP, a novel training-free compression framework that replaces redundant layers with adaptive singular parameters, leveraging the low-rank structure within LLMs to preserve performance with minimal overhead.
  • We introduce a gradient-based singular value selection mechanism, enabling efficient identification of critical components without relying on magnitude-based heuristics.
  • We conduct extensive experiments across ten datasets and five models from two major LLM families (LLaMA and Mistral), demonstrating that GRASP consistently achieves strong performance under both training-free and low-resource fine-tuning settings.
2 Method
Figure 1 illustrates the workflow of GRASP. The method consists of two main steps: Identifying redundant layers to be pruned (Sec 2.1) and replacing redundant layers with critical singular components guided by gradient-based attribution (Sec 2.2). Below, we describe each step in detail.
2.1 Redundant Layer Selection
The first step in GRASP is identifying redundant layers. These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance. Following prior works (Song et al., 2024; Chen et al., 2024), we use cosine similarity to quantify the degree of transformation in a given layer.
For a transformer layer with input hidden state $H_{i}\in\mathbb{R}^{d}$ and output hidden state $H_{i+1}\in\mathbb{R}^{d}$, the cosine similarity is computed as:
$\text{cos}(H_{i},H_{i+1})=\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$ (1)
A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant. Instead of directly removing these layers, GRASP compresses weight matrices with a gradient-based approach to retain critical internal transformations.
2.2 Layer Replacement with Adaptive Singular Parameters
Motivation.
GRASP operates under the hypothesis that redundant layers exhibit an inherent low-rank structure, allowing their functionality to be effectively approximated using low-rank matrices. Based on this insight, a straightforward approach is to apply singular value decomposition (SVD) to these layers. However, prioritizing components solely based on singular value magnitude does not necessarily correlate with downstream task performance (Hsu et al., 2022; Hua et al., 2025). To validate this point, we selectively zero out groups of singular values in the weight matrices of large language models and measure their impact on downstream tasks. As shown in Figure 2, we make two key observations: (1) The contribution of a singular value to downstream task performance is not solely determined by its magnitude; smaller values can be crucial for task performance in some cases. (2) Redundant layers exhibit a highly low-rank structure, where only a few singular directions dominate the model performance.
Figure 2: Sensitivity analysis of grouped singular value truncation. While singular values are typically ordered by magnitude, their impact on downstream performance does not follow the same order.
Key Design.
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone. Formally, the singular value decomposition of a weight matrix $W\in\mathbb{R}^{m\times n}$ is given by $W=U\Sigma V^{\top}$. Equivalently, $W$ can be represented as the sum of its rank-one components:
$W=\sum_{k=1}^{l}u_{k}\sigma_{k}v_{k}^{\top},$ (2)
where $l=\min(m,n)$, $\sigma_{k}$ denotes the $k$-th singular value, and $u_{k}$, $v_{k}$ are the $k$-th column vectors of the orthogonal matrices $U$ and $V$, respectively. For brevity, we use $\Phi_{k}$ to represent the singular group $\{u_{k},\sigma_{k},v_{k}^{\top}\}$. GRASP estimates the importance of each $\Phi_{k}$ using a small, general-purpose calibration dataset (e.g., WikiText-2), computing a sensitivity-based score that reflects its effect on model performance, given by:
$I(\Phi_{k})=T(\sigma_{k})+\sum_{i=1}^{m}T(u_{k,i})+\sum_{i=1}^{n}T(v_{k,i})$ (3)
where $T(\cdot)$ denotes the estimated change in loss when a certain parameter $\theta$ is zeroed out. This can be approximated by the second-order Taylor series expansion (LeCun et al., 1989), which is defined as:
$T(\theta)=\left|\theta^{\top}\nabla_{\theta}\mathcal{L}+\frac{1}{2}\theta^{\top}H\theta+\mathcal{O}(\|\theta\|^{3})\right|$ (4)
$\mathcal{L}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (5)
where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ with respect to the parameter $\theta$, and $H$ represents the corresponding Hessian matrix. To reduce computational overhead, we omit the second-order term and approximate the importance using only the first-order derivative (Hua et al., 2025, Kim et al., 2024):
$I(\Phi_{k})=\left|\sigma_{k}\frac{\partial\mathcal{L}}{\partial\sigma_{k}}\right|+\sum_{i=1}^{m}\left|u_{k,i}\frac{\partial\mathcal{L}}{\partial u_{k,i}}\right|+\sum_{i=1}^{n}\left|v_{k,i}\frac{\partial\mathcal{L}}{\partial v_{k,i}}\right|$ (6)
The first term in Eq. 6 captures the gradient projection of the loss $\mathcal{L}$ with respect to the singular value $\sigma_{i}$. The detailed proof is provided in Appendix A.1. By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values. A larger value indicates a greater importance of the singular group. This gradient-oriented attribution moves beyond heuristic magnitude-based criteria, enabling performance-aligned importance evaluation. Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-$k$ singular values most critical to model performance based on Eq. 6, and use them to replace the corresponding redundant layer.
2.3 Detailed Implementation of GRASP
GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network. We summarize the detailed algorithm in Algorithm 1.
Algorithm 1 GRASP: Gradient-based Retention of Adaptive Singular Parameters
1:Model $M$, Calibration set $D$, Retain ratio $r$
2:Compressed model $\tilde{M}$
3:Step1: Redundant Layer Selection:
4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers
5:Select top-$L$ layers with highest similarity as redundant
6:Step 2: Gradient-Guided Compression:
7:for each redundant layer $l$ (in reverse order) do
[…]
9:         SVD: $W=U\Sigma V^{\top}$
10:         Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq. 6
11:         Keep top-$r\%$ singular groups, reconstruct $\tilde{W}$
12:     end for
13:end for
14:return $\tilde{M}$
3 Experiments
In this section, we conduct comprehensive experiments to evaluate GRASP from three key perspectives. (1) We first compare our method with existing pruning-based LLM compression approaches to demonstrate its effectiveness (Section 3.2). (2) Next, we analyze the inference speed-up achieved by GRASP (Section 3.3). (3) Finally, we investigate the factors influencing our approach by performing ablation studies on the choice of calibration datasets and pruning strategies. (Section 3.4)
3.1 Experimental Setup
Below we detail the models, benchmarks, baselines and implementation details used in our experiments, with more experimental setups provided in Appendix A.2.
Models.
We evaluate GRASP on a range of large language models (LLMs) from two model families: the LLaMA family, including LLaMA-7B (Touvron et al., 2023a), LLaMA 2-7B, LLaMA 2-13B (Touvron et al., 2023b), and LLaMA 3.1-8B-Instruct (Dubey et al., 2024), as well as Mistral-7B (Jiang et al., 2023) from the Mistral family.
Baselines.
We compare GRASP against 8 structured pruning methods to substantiate its efficacy:
  • Layer-pruning methods We consider three representative layer-pruning methods as baselines: ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024) and LLM-Streamline (Chen et al., 2024).
  • Module-pruning methods We also select LLM-Pruner (Ma et al., 2023) and SliceGPT (Ashkboos et al., 2024) which prune the redundant modules in LLMs.
  • Low-rank Pruning methods Considering our method involves Gradient-based SVD, we also compare with other low-rank pruning methods: FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024).
We provide a detailed comparison of these pruning-based LLM compression methods in Appendix A.3.
Implementation Details.
To ensure a fair comparison, we randomly sample 512 data points from the WikiText-2 dataset as the calibration dataset. All experiments are conducted on NVIDIA A100-SXM4 (80GB) GPUs. Further experimental details can be found in Appendix A.2.
Methods OpenbookQA ARC_e WinoGrande HellaSwag ARC_c PIQA MathQA Average Percentage
Dense 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60 100.0%
LaCo 0.26 0.49 0.65 0.33 0.30 0.65 0.30 0.42 70.9%
ShortGPT 0.21 0.57 0.66 0.42 0.32 0.67 0.26 0.44 74.1%
SliceGPT 0.15 0.43 0.51 0.30 0.23 0.58 0.22 0.35 57.7%
GRASP 0.22 0.60 0.70 0.44 0.37 0.69 0.28 0.47 78.6%
Table 1: Zero-shot performance of GRASP and structured pruning baselines without post-training under a 20% compression ratio. Results are reported on seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
Method C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU Avg. Per.
Dense 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 49.2 100.0%
LLMPruner* 29.7 33.4 28.4 58.7 40.4 48.5 54.6 72.0 22.9 22 25.3 25.0 38.4 78.1%
SliceGPT* 31.5 31.6 18.5 59.9 43.3 49.6 47.5 68.3 27.0 29.4 28.8 24.8 38.4 78.0%
LaCo* 39.7 34.4 36.1 64.1 40.4 45.7 55.7 69.8 23.6 22.6 26.5 25.2 40.3 82.0%
ShortGPT* 40.2 34.4 21.5 67.3 40.4 51.7 59.7 69.0 35.2 34.7 44.6 28.9 44.0 89.4%
LLM-Streamline-FFN* 40.7 33.0 22.8 65.9 38.5 60.6 61.2 71.2 38.0 38.7 47.0 31.7 45.8 93.1%
LLM-Streamline-Layer* 43.3 33.0 24.1 67.5 36.5 59.2 61.1 71.5 34.8 37.0 45.5 29.4 45.2 92.0%
GRASP 44.6 35.1 26.2 68.4 41.4 63.2 62.7 73.3 35.1 36.1 43.1 30.7 46.7 94.9%
Table 2: Comparison between GRASP and structured pruning baselines with post-training compensation under a 20% compression ratio. Results marked with * are reported from (Chen et al., 2024). Bold values indicate the best performance.
Ratio Method Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA GSM8K Average Percentage
0% Original 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.09 0.46 100.0
20% FWSVD 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.00 0.28 60.8
ASVD 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.04 0.38 82.6
SVD-LLM 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.05 0.39 84.7
Ours 0.22 0.52 0.64 0.43 0.32 0.70 0.24 0.03 0.39 84.7
Ours* 0.24 0.59 0.63 0.5 0.35 0.73 0.25 0.04 0.42 91.3
30% FWSVD 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.00 0.26 56.5
ASVD 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.00 0.33 71.7
SVD-LLM 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.03 0.35 76.1
Ours 0.19 0.42 0.62 0.39 0.28 0.64 0.23 0.02 0.35 76.1
Ours* 0.24 0.54 0.64 0.46 0.32 0.69 0.24 0.04 0.40 87.0
40% FWSVD 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.00 0.27 58.7
ASVD 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.00 0.26 56.5
SVD-LLM 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.02 0.33 71.7
Ours 0.18 0.37 0.57 0.35 0.27 0.61 0.21 0.01 0.32 69.6
Ours* 0.22 0.49 0.63 0.43 0.3 0.68 0.23 0.02 0.38 82.6
50% FWSVD 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.00 0.26 56.5
ASVD 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.00 0.26 56.5
SVD-LLM 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.00 0.29 63.0
Ours 0.13 0.29 0.53 0.28 0.23 0.53 0.20 0.01 0.28 60.9
Ours* 0.18 0.4 0.56 0.35 0.26 0.61 0.21 0.02 0.32 69.6
[…]
3.2 Comparison with Pruning-based LLM Compression Methods
3.2.1 Experiment 1
Models and Benchmarks.
In this experiment, we evaluate GRASP against representative structured pruning baselines using a more modern LLM architecture—LLaMA 3.1-8B-Instruct—without applying any post-training compensation. The model is compressed to 20% of its original size and evaluated on seven commonsense reasoning benchmarks, including WinoGrande (Sakaguchi et al., 2020), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-e, ARC-c (Clark et al., 2018), and MathQA (Amini et al., 2019). All tasks are tested in a zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024).
Main Results.
As shown in Table 1, GRASP achieves the highest average accuracy across seven commonsense reasoning benchmarks, consistently outperforming all baseline methods. In particular, GRASP improves over SliceGPT by 34% in average accuracy and outperforms LaCo by 12%. To further evaluate the generalizability of GRASP across different LLM architectures, we additionally conduct experiments on LLaMA 2-7B, LLaMA 2-13B, and Mistral-7B. Detailed results are presented in Appendix A.4. Notably, GRASP demonstrates significantly improved robustness across models, effectively mitigating the variability in pruning sensitivity across diverse architectures.
3.2.2 Experiment 2
Models and Benchmarks.
In this section, following prior research, we compress the LLaMA 2-7B model under a 25% compression ratio and evaluate the compressed model on a broad set of natural language understanding (NLU) and question-answering (QA) benchmarks, including CMNLI (Xu et al., 2020), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), CHID (Zheng et al., 2019), WSC (Levesque et al., 2012), CommonsenseQA (Talmor et al., 2018), BoolQ (Clark et al., 2019), MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2023), Race (Lai et al., 2017) and C3 (Sun et al., 2020). We utilized the OpenCompass evaluation framework (Contributors, 2023) and report accuracy as the evaluation metric for all benchmarks under the PPL mode, following the same evaluation protocol as LaCo (Yang et al., 2024).
Main Results.
To ensure a fair comparison, we constrain the number of trainable parameters to remain approximately the same across all methods by retaining only 10% of the parameters in each redundant layer and allowing only these parameters to be trainable. Notably, for GRASP’s post-training compensation process, we fine-tune the compressed model on Alpaca (Taori et al., 2023) for only one epoch to guarantee efficiency. Additional implementation details are provided in Appendix A.2. As shown in Table 2, GRASP consistently outperforms the best-performing baselines LLM-Streamline on average and achieves a 94.9% of the original model performance at a 25% compression ratio. In addition to achieving superior accuracy, GRASP also demonstrates more stable and faster convergence during post-training, which we attribute to its preservation of critical singular components.
3.2.3 Experiment 3
Models and Benchmarks.
In this section, we further compare our method against state-of-the-art structured low-rank pruning approaches—FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024) on the LLaMA-7B model under various compression ratio. 8 datasets are used as evaluation benchmarks including seven commonsense reasoning datasets as Experiment 1 and one natural language generation (NLG) benchmark GSM8K (Cobbe et al., 2021). For all benchmarks, we report the zero-shot accuracy as the evaluation metric.
Main Results.
To evaluate the performance and stability of our proposed method, we conduct experiments under various compression ratios ranging from 20% to 50%. Table 3 summarizes the results for different methods. The results demonstrate that our proposed GRASP consistently outperforms the baseline methods on most benchmarks. Specifically, GRASP retains more than 91% of the original performance at a 20% compression ratio and 87% under the compression ratio of 30%. More importantly, with fast and resource-efficient post-training compensation, GRASP enables rapid accuracy recovery, achieving 70% of the original model performance even at a 50% compression ratio.
Furthermore, to assess the generalizability of GRASP across different LLM architectures, we compare its performance against structured low-rank pruning baselines under a 20% compression ratio on four models from two distinct LLM families: LLaMA 2-7B, LLaMA 2-13B, LLaMA 3.1-8B-Instruct, and Mistral-7B. As shown in Figure 3, GRASP consistently outperforms all baseline methods across architectures and exhibits greater robustness across different model families. The only exception is on LLaMA 2-13B, where GRASP slightly underperforms SVD-LLM. However, this performance gap can be quickly recovered through lightweight compensation. The detailed results are provided in Appendix A.5.
Figure 3: Comparison between our method and low-rank pruning baselines on four different LLMs. Average accuracy is reported across seven commonsense reasoning benchmarks: OpenBookQA, WinoGrande, HellaSwag, ARC-easy, ARC-challenge, PIQA, and MathQA.
Figure 4: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different calibration datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
3.3 Compression Costs and Inference Efficiency of GRASP
GRASP enables low-cost compression of LLMs while improving inference efficiency on real hardware. To evaluate its acceleration benefits, we measure the throughput (tokens per second) of the original LLaMA2-7B and its GRASP-compressed counterpart under varying sequence lengths and batch sizes. As shown in Figure 5, GRASP consistently improves generation speed and achieves acceleration comparable to direct layer removal. Notably, although GRASP retains a small subset of parameters within redundant layers to mitigate the performance drop caused by layer removal, these retained components are extremely low-rank and incur negligible inference overhead while preserving task performance. Additionally, we measure the compression time of GRASP and other structured pruning baselines when compressing LLaMA2-7B on an NVIDIA A100 GPU under a 25% compression ratio. As shown in Table 4, GRASP is able to compress the model efficiently while maintaining strong downstream performance, demonstrating its practicality for real-world deployment.
Model Pruning Time (h) Compensation Time (h)
LaCO 0.05 1.2
SliceGPT 0.6 0.76
LLM-Streamline 0.03 0.7
GRASP 0.16 ×
GRASP* 0.16 0.66
Table 4: Compression time of GRASP and structured pruning baselines on LLaMA2-7B under a 25% compression ratio on a single A100 GPU. “×” indicates that GRASP does not require post-training compensation.
Figure 5: Throughput of LLaMA2-7B and GRASP compressed model under 25% compression ratio on a single A100 GPU. Top: Throughput across different sequence lengths (batch size = 32). Bottom: Throughput across different batch sizes (sequence length = 32).
3.4 Ablation Study
Calibration Data.
We conduct ablation studies to assess the sensitivity of GRASP to the choice and size of the calibration dataset used for gradient-based singular value attribution. Table 5 and Figure 4 present the results on LLaMA3.1-8B-Instruct when varying the calibration dataset and the number of samples drawn from WikiText-2. The results indicate that GRASP remains robust across different calibration dataset choices and performs reliably even with limited calibration data. Remarkably, the method achieves strong performance with as few as 64 samples, demonstrating its effectiveness in low-data regimes.
Ablation Type
Calibration
Dataset
WikiText-2 PTB
Average
Accuracy
Varying Dataset WikiText-2 37.86 63.97 47.12
C4 40.54 71.42 46.17
Varying Number 64 46.5 86.51 47.06
128 39.91 76.41 46.93
256 38.73 79.13 46.67
512 37.86 63.97 47.12
Table 5: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct at a 20% compression ratio. Results are reported on the WikiText-2, PTB datasets (measured by $\downarrow$) and the average accuracy across seven commonsense reasoning benchmarks: OpenBookQA, WinoGrande, HellaSwag, ARC-easy, ARC-challenge, PIQA, and MathQA.
One-shot Pruning vs Iterative Pruning
As detailed in Section 2.3, GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. We present the ablation results in Table 10 in Appendix A.8, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
4 Discussion
The effectiveness of GRASP stems from two key design principles that distinguish it from conventional structured pruning methods.
Leveraging Low-Rank Redundancy.
GRASP builds on the observation that the redundant layers in LLMs exhibit inherent low-rank characteristics. Rather than removing these layers entirely, which can lead to information loss and degraded performance, GRASP retains a small subset of critical singular directions identified through a gradient-based attribution mechanism. This approach allows the model to retain essential functional capacity while discarding non-informative parameters, leading to efficient yet accurate compression.
Preserving Informative Subspaces for Fast Recovery.
By maintaining only the most influential parameters within redundant layers, GRASP avoids the need for costly retraining while facilitating faster convergence when post-training compensation is applied. This contrasts with approaches like LLM-Streamline (Chen et al., 2024) that insert randomly initialized or dense lightweight modules, which require more data and training time to approximate the original function. GRASP’s retention of critical low-rank components allows it to preserve the spectral alignment of the original model (Oymak et al., 2019, Kamalakara et al., 2022, Shuttleworth et al., 2024), resulting in more stable optimization dynamics and enhanced sample efficiency during post-training compensation.
5 Related Work
Structured pruning aims to reduce the size and computational cost of large language models (LLMs) by removing entire components such as layers, neurons, or dimensions, while maintaining model performance. Among this, layer pruning is a kind of structured pruning technique that eliminates redundant layers within Large Language Models (LLMs). Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers. SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs. LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
Although effective in reducing model size, layer pruning disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al. (2024). To mitigate this, post-training compensation methods have been proposed. Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities. However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024). LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers. While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
Another line of structured pruning research focuses on low-rank approximation, where Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss. Recent methods have enhanced SVD to reduce LLM compression error. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices. In contrast to these approaches, GRASP integrates gradient-based attribution into the low-rank decomposition process, enabling both efficient compression and critical parameter retention.
6 Conclusion
In this work, we proposed GRASP, a novel compression framework that replaces redundant layers in large language models with a small set of adaptively selected singular parameters. By leveraging the low-rank structure of redundant layers and incorporating gradient-based attribution, GRASP identifies critical components that preserve model functionality with minimal parameter overhead. It operates in a training-free manner and enables efficient post-training recovery with limited data. Extensive experiments across diverse LLM architectures and benchmarks demonstrate that GRASP consistently outperforms existing structured pruning methods in both accuracy and efficiency.
7 Limitations
While GRASP achieves competitive performance and compression efficiency, it also has several limitations that merit further exploration.
First, our method relies on the assumption that layer redundancy can be identified via output similarity (e.g., cosine similarity between hidden states). While effective in practice, this heuristic may overlook more nuanced forms of redundancy that arise from distributed or task-specific behaviors within deeper model layers.
Second, GRASP depends on access to gradient information and a small calibration dataset to compute attribution scores. Although the data requirement is minimal and the method is training-free in its core form, this may limit applicability in strictly black-box or privacy-sensitive settings where gradients or internal representations are inaccessible.
We also note that our experiments are conducted on models up to 13B parameters and primarily in English-language tasks. Extending GRASP to multilingual or much larger-scale models is a promising direction for future work, especially as the scale and diversity of LLMs continue to grow.
Ethical Considerations
Our research adheres to the ACL Code of Ethics, ensuring transparency, responsible use of data, and consideration of potential social impacts. All datasets used in this work are publicly available and have been appropriately cited, ensuring compliance with data usage agreements and privacy regulations.
While GRASP is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight. To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
Moreover, we acknowledge the computational resources required for training and testing large language models. To minimize environmental impact, we conducted experiments on energy-efficient hardware (NVIDIA A100 GPUs) and report our computational cost transparently. Further details can be found in the Appendix.
Appendix A Appendix
A.1 The gradient of singular values
For a weight matrix $W\in\mathbb{R}^{m\times n}$ in the selected redundant layers, its differential form can be expressed as:
$\partial W=\partial U\Sigma V^{T}+U\partial\Sigma V^{T}+U\Sigma\partial V^{T}$
$U^{T}\partial WV=U^{T}\partial U\Sigma+\partial\Sigma+\Sigma V^{T}\partial V$
Since both $U$ and $V$ are orthogonal matrices, we have:
$U^{T}U=I_{m},\quad V^{T}V=I_{n}$
$\partial U^{T}U+U^{T}\partial U=O_{m},\quad\partial V^{T}V+V^{T}\partial V=O_{n}$
This implies that $U^{T}dU$ and $dV^{T}V$ are asymmetric matrices. Therefore, the diagonal elements of $U^{T}dU\Sigma$ and $\Sigma V^{T}dV$ are zero, leading to the diagonal elements of $U^{T}\partial WV$ being:
$I_{k}\odot U^{T}\partial WV=\partial\Sigma$
where $I_{k}$ represents the $k\times k$ identity matrix, $\odot$ denotes element-wise multiplication.
For a singular value $\sigma_{i}$, its differential form can be written as:
$\partial{\sigma_{i}}=u_{i}^{T}\partial Wv_{i}$
Since $\sigma_{i}$ is a scalar, we have:
$\partial\sigma_{i}$ $=\text{tr}(\partial\sigma_{i})$
$=\text{tr}(u_{i}^{T}\partial Wv_{i})$
$=\text{tr}[(u_{i}v_{i}^{T})^{T}\partial W]$
thereby, the derivative of $\sigma_{i}$ with respect to W is:
$\frac{\partial\sigma_{i}}{\partial W}=u_{i}v_{i}^{T}$
For a calibration dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
$\frac{\partial L}{\partial\sigma_{i}}=u_{i}^{T}\frac{\partial L}{\partial W}v_{i}$
Then, for all the singular values $\Sigma$, we have:
$\frac{\partial L}{\partial\Sigma}=I_{k}\odot U^{T}\left(\frac{\partial L}{\partial W}\right)V$
A.2 Experimental Setup and Hyperparameters Configuration
To ensure a fair comparison, all experimental setups are consistent across all methods. In the following, we describe the experimental setup and hyperparameters configuration in detail.
Hyperparameters Configurations
For post-training compensation, all models compressed by GRASP are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$. All our experiments are conducted on a single A100 GPU with mixed precision enabled. Table 6 provides the detailed configurations of post-training compensation.
HyperParameters Setting
Dataset Alpaca
Huggingface Dataset Path yahma/alpaca-cleaned
Batch Size 32
Micro Batch Size 4
Epochs 1
Learning Rate 3.00E-04
Max Length 256
Train on Inputs TRUE
Add EOS Token FALSE
LoRA-Rank 256
LoRA-Alpha 16
LoRA-Dropout 0.05
LoRA-Target-Modules
q_proj, k_proj, v_proj, o_proj,
up_proj, down_proj, gate_proj
Prompt-Template Alpaca Template
Table 6: Experimental setup and hyperparameters configurations.
Method Metric Calibration Data Need Post-Training Training Data Training Dataset Size Training Module
ShortGPT Cosine Similarity WikiText-2 No None None None
LaCo Cosine Similarity WikiText-2 Optional Unpublished 1B Full Parameters
SliceGPT PCA
WikiText-2
Alpaca
Optional Alpaca 5k Full Parameters
Shortened LLaMA
Taylor
Perplexity
BookCorpus Optional
SlimPajama
Alpaca
627B
50k
Full Parameters
LoRA-Adapter
LLM-Streamline Cosine Similarity WikiText-2 Yes SlimPajama 30k
Lightweight
Network
GRASP Cosine Similarity WikiText-2 Optional Alpaca 50k
Low-rank
Modules
Table 7: Comparison of pruning-based LLM compression methods, where the metric indicates the criterion used to identify redundant modules. "Optional" refers to methods that can either work without post-training or recover performance through post-training. Shortened LLaMA consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
A.3 Comparison of Concurrent Structured Pruning Methods
We provide a detailed comparison of concurrent structured pruning LLM compression methods, and the results are summarized in Table 7.
A.4 Detailed Results on Commonsense Reasoning Benchmarks
In this section, we provide detailed results for GRASP and baseline methods on the commonsense reasoning benchmarks using LLaMA2-7B, LLaMA2-13B, and Mistral-7B. These results extend the main experiments presented in Table 1, offering a comprehensive view of GRASP’s performance across different model architectures.
Table 8 reports accuracy without post-training compensation. The results demonstrate that GRASP consistently maintains strong accuracy and demonstrates robustness across model scales and families.
A.5 More Results on Other Models
To further validate the generalizability of GRASP across diverse LLM architectures, we present additional results under a 20% compression ratio on three representative models beyond LLaMA-7B, LLaMA2-13B, LLaMA3.1-8B-Instruct, and Mistral-7B. Table 9 summarizes the performance comparison between GRASP and structured low-rank pruning baselines, including FWSVD, ASVD, and SVD-LLM, across eight evaluation benchmarks.
Consistent with our findings in the main text, GRASP achieves superior or comparable accuracy across most benchmarks. While SVD-LLM slightly outperforms GRASP on LLaMA2-13B, our method demonstrates stronger robustness across model families and benefits from more stable post-compensation recovery. These results highlight GRASP’s effectiveness as a generalizable and architecture-agnostic compression strategy.
Model Method Ratio Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
Mistral-7B-v0.1 Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
LaCo 21.1% 0.20 0.35 0.58 0.26 0.25 0.53 0.24 0.34
ShortGPT 21.1% 0.19 0.57 0.68 0.46 0.37 0.71 0.26 0.46
SliceGPT 20.0% 0.19 0.51 0.59 0.35 0.25 0.61 0.23 0.39
GRASP 20.0% 0.21 0.56 0.68 0.43 0.38 0.67 0.26 0.46
LLaMA2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
LaCo 18.1% 0.26 0.48 0.59 0.42 0.32 0.69 0.24 0.43
ShortGPT 21.1% 0.23 0.49 0.63 0.42 0.31 0.68 0.23 0.43
SliceGPT 21.5% 0.22 0.54 0.61 0.37 0.28 0.63 0.23 0.41
GRASP 21.6% 0.24 0.54 0.63 0.43 0.33 0.71 0.23 0.44
LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
LaCo 19.5% 0.28 0.52 0.63 0.43 0.33 0.70 0.25 0.45
ShortGPT 22.1% 0.24 0.50 0.63 0.46 0.33 0.7 0.24 0.44
SliceGPT 20.0% 0.29 0.59 0.65 0.39 0.32 0.64 0.24 0.45
GRASP 20.0% 0.26 0.61 0.66 0.47 0.35 0.73 0.24 0.47
LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
LaCo 19.0% 0.26 0.49 0.65 0.33 0.30 0.65 0.30 0.42
ShortGPT 21.7% 0.21 0.57 0.66 0.42 0.32 0.67 0.26 0.44
SliceGPT 20.0% 0.15 0.43 0.51 0.30 0.23 0.58 0.22 0.35
GRASP 20.0% 0.22 0.60 0.70 0.44 0.37 0.69 0.28 0.47
Table 8: Zero-shot performance of GRASP and pruning-based without post-training baselines under 20% compression ratio. Results are reported on seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
Original 5.25 0.59 5.68 0.52 5.47 0.53 7.21 0.60
FWSVD 6357 0.32 1727 0.32 2360 0.31 3256.7 0.29
ASVD 19.28 0.4 11.14 0.44 9.70 0.46 2443.99 0.30
SVD-LLM 10.21 0.41 7.94 0.44 8.50 0.47 - -
Ours 18.42 0.45 14.79 0.44 16.12 0.44 37.86 0.47
Ours* 11.62 0.51 10.19 0.47 9.59 0.48 14.13 0.53
Table 9: Perplexity($\downarrow$) of GRASP and low-rank pruning baselines on the WikiText-2 datasets and the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio. "†" indicates that we refer to the results in the original paper. The best performance is marked in bold.
A.6 Evaluation Results on LongBench
In this section, we present the detailed results of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench, which are presented in Table 11. The results illustrate that GRASP with post-training compensation still maintains superior performance on long-form reasoning and complex generative tasks.
A.7 Robustness of GRASP towards different calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 4, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of data. Tables 12 and 13 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
A.8 Robustness of GRASP towards Different Pruning Strategy
In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 10
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22
Table 10: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
Model Summarization Few-shot Learning Synthetic Task
1-1 1-2 1-3 1-4 1-1 2-1 2-2 2-3 2-4 3-1 3-2
LLaMA3.1-8B-Instruct 28.35 20.04 25.85 14.51 63 51.3 39.29 16.5 1.64 9.67 17.17
GRASP 25.76 19.41 25.97 8.99 59.5 67.44 38.41 18 1 10 18
Model One-Doc QA Multi-Doc QA Code Completion Average
4-1 4-2 4-3 4-4 5-1 5-2 5-3 5-4 6-1 6-2 ALL
LLaMA3.1-8B-Instruct 9.96 6.04 18.75 14.27 11.27 11.57 7.65 34.16 56.47 51.34 26.09
GRASP 15.32 20.47 34.52 16.46 31.33 26.18 10.28 29.86 33.74 38.09 26.26
Table 11: Performance comparison of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench. The datasets are grouped as follows: (1-1 to 1-4) denote GovReport, QMSum, MultiNews, and VCSUM; (2-1 to 2-4) denote TREC, TriviaQA, SAMSum, and LSHT; (3-1 to 3-3) denote PassageCount, PassageRetrieval-en, and PassageRetrieval-zh; (4-1 to 4-4) denote NarrativeQA, Qasper, MultiFieldQA-en, and MultiFieldQA-zh; (5-1 to 5-4) denote HotpotQA, 2WikiMultihopQA, MuSiQue, and DuReader; (6-1 to 6-2) denote LCC and RepoBench-P.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
WikiText-2 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
C4 40.54 71.42 24 57.91 67.72 42.11 36.69 67.25 27.5 46.17
Table 12: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as calibration datasets.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
WikiText-2-64 46.5 86.51 22.6 59.97 69.3 44.36 37.29 68.5 27.37 47.06
WikiText-2-128 39.91 76.41 22.8 60.23 69.93 44.24 35.92 67.9 27.5 46.93
WikiText-2-256 38.73 79.13 21.8 59.89 70.24 44.23 36.26 67.19 27.07 46.67
WikiText-2-512 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 13: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.